### v1

# Secret-Key Generation from Private Identifiers under Channel Uncertainty

Abstract
This study investigates secret-key generation for device authentication using physical identifiers, such as responses from physical unclonable functions (PUFs). The system includes two legitimate terminals (encoder and decoder) and an eavesdropper (Eve), each with access to different measurements of the identifier. From the device identifier, the encoder generates a secret key, which is securely stored in a private database, along with helper data that is saved in a public database accessible by the decoder for key reconstruction. Eve, who also has access to the public database, may use both her own measurements and the helper data to attempt to estimate the secret key and identifier. Our setup focuses on authentication scenarios where channel statistics are uncertain, with the involved parties employing multiple antennas to enhance signal reception. Our contributions include deriving inner and outer bounds on the optimal trade-off among secret-key, storage, and privacy-leakage rates for general discrete sources, and showing that these bounds are tight for Gaussian sources.
Index Terms:
Capacity region, compound channels, multiple outputs, key generation, privacy leakage, PUFs.
I Introduction
The Internet of Things (IoT) is a rapidly growing technology that enables numerous sensors and small-chip devices to interact and exchange information over the internet. However, ensuring security and privacy in IoT communications presents significant challenges compared to conventional networks due to the diverse range of applications and resource constraints of these devices [2]. To help address these difficulties, recent efforts have focused on developing security protocols at the physical layer for authenticating devices.
Secret-key generation using physical identifiers, such as responses from unclonable functions (PUFs), is a promising protocol for device authentication because it offers several advantages, including simple designs, low costs, and eliminating the need to save the secret key on the device [3]. A PUF is defined as a physical function that for a given input (challenge), provides an output (response) that serves as a unique identifier for each device [4, 5]. Some examples of PUFs are static random access memory (SRAM) PUFs and ring-oscillator (RO) PUFs.
Figure 1: An authentication scheme based on secret-key generation with PUFs. The eavesdropper (Eve) is a passive adversary who is interested in learning the secret key and the source identifier, but does not interfere with the communication mechanism of the system.
Secret-key authentication using PUFs is illustrated in Fig. 1 [6]111[6] does not consider the presence of Eve, but is included in the figure to facilitate understanding of our system model in Section II. and consists of an enrollment phase and an authentication phase. During the enrollment phase, the terminal (encoder) challenges the device, i.e., the PUF embedded in the device, with a challenge C and gets a response $\widetilde{\rm R}$, from which the encoder generates a secret key and helper data. The secret key is securely stored in a private database, while the helper data are saved in a public database, which can be accessed by both the decoder and Eve. In the authentication phase, the terminal (decoder) challenges the device by sending the same challenge C, which produces a different response R due to noise effects. The decoder reconstructs the secret key based on both the response and helper data from the public database and then compares it with the one saved in the private database. If they match, the device is successfully authenticated; otherwise, the authentication fails.
In subsequent discussions, the response of a PUF unaffected by noise is referred to as the source identifier. The responses of a PUF observed at the terminals and Eve through communication channels are called the observed identifiers.
I-A Motivations
We study the capacity region of secret-key generation from source identifiers for a setup involving compound authentication channels with multiple outputs. The motivation for considering compound-channel settings is to capture a situation in which device authentication takes place in environments where the channel statistics may not be perfectly known. This contrasts with most previous studies, which assume that the encoder and decoder have perfect knowledge of the source and channel statistics of the systems.
For example, as shown in Fig. 1, consider a situation where the decoder needs to authenticate a flying drone. As the channel state information (CSI) of the channel from the drone to the decoder may fluctuate, it makes it difficult for the decoder to obtain the exact CSI. Additionally, Eve is unlikely to share her CSI with the legitimate terminals. Thus, compound channels are used to model the channels to the decoder and Eve. In this setting, the encoder and decoder do not possess precise CSI of the relevant channels but are aware that these channels belong to certain predefined sets.
Additionally, we consider multiple outputs for the channels to the encoder, decoder, and Eve to capture the circumstance where these parties may deploy multiple antennas to enhance signal reception. Note that having more antennas can increase the correlation between Eve’s observation and the source identifier, giving her an advantage in learning the secret key and the source identifier. Hence, in this setting, we want to quantify the potential leakage to Eve from both security and privacy perspectives.
Finally, in practice, certain types of PUFs produce continuous-value identifiers. For example, the source of RO PUFs can be modeled as a Gaussian distribution [7]. Additionally, a number of communication channels are sometimes approximated as additive white Gaussian noise (AWGN) channels. This motivates us to study setups with Gaussian sources and AWGN channels.
I-B Related Work
Secret-key generation using PUFs [^2] PUF and biometric identifiers share similar characteristics, and thus, the theoretical results developed for one can be applied to the other as well [3]. has been studied from information-theoretic perspectives in [8, 9]. Later, several extensions of this model were found in [10, Ch. 4] for Gaussian sources, [11] for separated and combined enrollments, and [12] for multiple rounds of enrollments and authentications. Limited storage rate was introduced to the model in [13]. Furthermore, the fundamental limits among secret-key, storage, and privacy-leakage rates when Eve also has a correlated sequence of the source were characterized in [14] for discrete sources and [15] for Gaussian sources. Note that this model is similar to the key-generation problem [16], [17], [18], but an additional privacy constraint is imposed in the problem formulation to limit information leakage on the source identifier.
More recent works have considered a setup that incorporates a noisy channel in the enrollment phase [19, 20, 21]. The channel is modeled to account for the noise introduced to the source identifier during the enrollment process, providing a more general framework, as signals generated by a PUF are inherently affected by noise. Further progress in this setting has been investigated in [22, 23, 24, 25], addressing user identification.
Secret-key generation with PUFs for compound sources has been studied in [26, 27] for the generated-secret (GS) model and the chosen-secret (CS) model [8, Sect. III]. Some extensions on this setting are explored in [28, 29] to incorporate user identification. Similar problems can be found in [30, 31, 32, 33] for key generation where the privacy constraint is not imposed and in [34, 35, 36, 37] for compound wiretap channels.
I-C Main Challenges and Contributions
The setup in [26, 27] focuses on uncertainty of the source statistics, whereas our work addresses channel uncertainty, when the channels can have multiple states whose realizations are unknown to the terminals. Additionally, compared to these prior studies [26, 27], our model incorporates enrollment noise and side information at Eve. Moreover, while previous works [26, 27] strictly focus on discrete sources, our study extends the discussion to investigate the capacity regions for both discrete and Gaussian cases. Below, we highlight some technical challenges encountered during the analysis of our newly considered setting.
For general discrete sources, since the encoder and decoder do not know the true CSI, in the achievability proof, the secret key’s uniformity and secrecy-leakage constraints must be satisfied for all possible receiver-eavesdropper channel states. In point-to-point communications [20], these constraints can be handled by bounding conditional entropy for joint typical sequences. However, in our analysis, an additional step, showing that there exists a random variable that simultaneously achieves the maximum mutual information across all Eve’s channels and is jointly typical with the source random variables with high probability, is essential to analyze these constraints. Moreover, unlike [20], the obtained inner and outer bounds involve optimizations over the channel states. Specifically, the inner region first involves an optimization carried over the distributions of auxiliary random variables, and then a minimization of the index pair for channel states. In contrast, the order of these two operations is reversed in the outer region. This leads to a gap between the inner and outer bounds, similar to the conclusion drawn in [34] for compound wiretap channels.
However, we show that, for a noiseless enrollment channel, our inner and outer regions coincide for Gaussian sources, providing a complete capacity characterization. The main challenging aspect arises in proving the converse part. Given the multiple-antenna settings at the legitimate terminals and Eve, the vector-form observations are not stochastically degraded in general [38]. We use sufficient statistics [39] to convert the vector problem into a scalar one. However, after this conversion, showing that all constraints of the original problem definition are preserved is challenging, and it is unclear whether the same expressions of the outer bounds for general discrete sources also hold for the scalar variables. Therefore, we cannot directly apply the technique in [40, Appx. B] to eliminate the second auxiliary random variable. In this paper, we instead derive new single-letter expressions of outer bounds for the Gaussian case using scalar random variables.
Another difficulty arises in proving the converse for the parametric expression of Gaussian sources. In the analysis of the model without side information at Eve [10, Appx. D], [41], the conditional entropy power inequality (EPI) plays an important role. However, the EPI is insufficient to prove the converse for all possible values of the optimization parameter in our problem. To overcome this issue, we adopt a distinct method introduced in [42, Sect. IV-C], using Fisher information. This approach enables us to derive the outer region that coincides with the inner one for any value of the optimization parameter.
Our main contributions are summarized as follows:
  • We derive inner and outer bounds on the capacity regions of secret-key, storage, and privacy-leakage rates of the GS and CS models for discrete sources.
  • We provide a complete characterization of the capacity regions of the GS and CS models for Gaussian sources by demonstrating the existence of a saddle point at which the inner and outer bounds coincide. Moreover, numerical calculations for the GS model are provided to illustrate how the change of the number of antennas at the decoder and Eve affects the secret-key and storage rates.
Our results recover, as special cases, results derived in previous works. For discrete sources, when only the channel to the decoder is compound and all channels have a single output, the inner and outer bounds match the preliminary result given in [1, Props. 1 and 2]. Additionally, the inner and outer bounds are tight for single-output and non-compound channels, and recover [20, Th. 3 and 4] without action cost. For Gaussian sources, as detailed in Section III-B, our results recover as special cases the capacity regions derived in [10, Ch. 4] and [21] for single-output and non-compound channels.
I-D Notation and Paper Organization
$\mathbb{R}_{+}$ is the set of non-negative real numbers. For any $a,b\in\mathbb{R}$, define $[a:b]\triangleq[\lfloor a\rfloor,\lceil b\rceil]\cap\mathbb{N}$. Italic uppercase $X$ and lowercase $x$ denote a random variable and its realization, respectively. Boldface letters $\bf X$ and $\bf x$ represent a collection of random variables and its realization. $X^{n}$ denotes the vector $(X_{1},{\dots},X_{n})$ and $X_{t}$ represents the $t$-th element in the vector. $X_{k}^{t}$ stands for a partial sequence $(X_{k},\dots,X_{t})$ for any $[k:t]\subseteq[1:n]$. $\sigma^{2}_{X}$ and ${\bf\Sigma}_{\mathbf{Y}}$ denote the variance of $X$ and the covariance matrix of ${\bf Y}$. $\mathcal{N}(0,\sigma^{2})$ denotes the Gaussian distribution with zero mean and variance $\sigma^{2}$. $\mathcal{T}^{n}_{\delta}(X)$ denotes the set of $\delta$-strongly typical sequences according to $P_{X}$ [43] and the random variable inside the parentheses is omitted, e.g., $\mathcal{T}^{n}_{\delta}$, when it is clear from the context. Additionally, the set of conditionally $\delta$-typical sequences is denoted as $\mathcal{T}^{n}_{\delta}(XY|z^{n})\triangleq\{(x^{n},y^{n}):(x^{n},y^{n},z^{n})\in\mathcal{T}^{n}_{\delta}\}$.
The remainder of the paper is organized as follows. In Section II, we state the problem definitions for the GS and CS models. We present our main results in Section III. The proofs of our main results are available in the appendices. Finally, we give concluding remarks in Section IV.
II Problem Statement
[…]
$P_{\tilde{\bf X}^{n}X^{n}{\bf Y}^{n}_{k}{\bf Z}^{n}_{l}}$ $\triangleq\textstyle\prod_{t=1}^{n}P_{\tilde{\bf X}_{t}|X_{t}}\cdot P_{X_{t}}\cdot P_{\mathbf{Y}_{k,t}\mathbf{Z}_{l,t}|X_{t}}.$ (1)
Secret-key generation strategies are formally defined below. Let $\mathcal{S}\triangleq[1:2^{nR_{S}}]$ and $\mathcal{J}\triangleq[1:2^{nR_{J}}]$.
Figure 2: Illustration of the system model in State $(k,l)$.
Definition 1 (GS model).
For the GS model, a $(2^{nR_{S}},n,R_{J},R_{L})$ secret-key generation strategy consists of:
  • Encoding mapping $e:\tilde{\mathbfcal{X}}^{n}\rightarrow\mathcal{J}\times\mathcal{S}$;
[…]
and operates as follows:
  • The encoder observes $\tilde{\mathbf{X}}^{n}$ and generates the helper data $J\in\mathcal{J}$ and a secret key $S\in\mathcal{S}$, as $(J,S)\triangleq e(\tilde{\mathbf{X}}^{n})$;
  • The helper data $J$ is stored in a public database, accessible to anyone;
  • From $\mathbf{Y}^{n}_{k}$ and $J$, retrieved from the public database, the decoder estimates $S$ as $\hat{S}_{k}\triangleq d(\mathbf{Y}^{n}_{k},J)$.
Definition 2 (CS model).
For the CS model, a $(2^{nR_{S}},n,R_{J},R_{L})$ secret-key generation strategy consists of:
  • A secret key $S$, chosen uniformly at random in $\mathcal{S}$ and independently of $(\tilde{\mathbf{X}}^{n},X^{n},\mathbf{Y}^{n}_{k},\mathbf{Z}^{n}_{l})$;
  • Encoding mapping $e:\tilde{\mathbfcal{X}}^{n}{}\times\mathcal{S}\rightarrow\mathcal{J}$;
[…]
and operates as follows:
  • From $\tilde{\mathbf{X}}^{n}$ and $S$, the encoder generates $J\triangleq e(\tilde{\mathbf{X}}^{n},S)$;
  • The helper data $J$ is saved in a public database, accessible to anyone;
  • From $\mathbf{Y}^{n}$ and $J$, the decoder estimates $S$ as $\hat{S}_{k}\triangleq d(\mathbf{Y}^{n}_{k},J)$.
In the following, we write $(\displaystyle\max_{k\in\mathcal{K}},\displaystyle\min_{k\in\mathcal{K}})$ and $(\displaystyle\max_{l\in\mathcal{L}},\displaystyle\min_{l\in\mathcal{L}})$ as $(\displaystyle\max_{k},\displaystyle\min_{k})$ and $(\displaystyle\max_{l},\displaystyle\min_{l})$, respectively, for simplicity.
Definition 3 (GS model).
A tuple of secret-key, storage, and privacy-leakage rates $(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}$ is achievable for the GS model if, for sufficiently small $\delta>0$ and large enough $n$, there exist pairs of encoders and decoders satisfying
$\max_{k}\mathbb{P}\{\widehat{S}_{k}\neq S\}$ $\leq\delta,$ (2)
$H(S)+n\delta$ $\geq\log|\mathcal{S}|\geq n(R_{S}-\delta),$ (3)
$\log{|\mathcal{J}|}$ $\leq n(R_{J}+\delta),$ (4)
$\max_{l}I(S;J,\mathbf{Z}^{n}_{l})$ $\leq n\delta,$ (5)
$\max_{l}I(X^{n};J|\mathbf{Z}^{n}_{l})$ $\leq n(R_{L}+\delta).$ (6)
$\mathcal{R}_{G}$ is defined as the closure of the set of all achievable rate tuples for the GS model, and it is called the capacity region.
Definition 4 (CS model).
A tuple of secret-key, storage, and privacy-leakage rates $(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}$ is achievable for the CS model if, for sufficiently small $\delta>0$ and large enough $n$, there exist pairs of encoders and decoders that satisfy all the conditions (2)–(6) with replacing (3) with $\log|\mathcal{S}|\geq n(R_{S}-\delta).$ Let $\mathcal{R}_{C}$ be the capacity region of the CS model.
In Definition 3, (2) denotes the reliability constraint, (3) is the uniformity requirement of the generated secret key, (4) is the storage rate constraint, (5) is the secrecy-leakage constraint, evaluating the information about the secret key leaked to Eve, and (6) is the privacy-leakage constraint, quantifying the amount of information leaked to Eve regarding the source identifier via the helper data given Eve’s side information.
III Main Results
In this section, we first present inner and outer bounds for the capacity region of the GS and CS models with discrete sources in Section III-A. We then present the capacity regions for Gaussian sources in Section III-B.
III-A Discrete Sources
Proposition 1 (Inner bounds).
We have
$\mathcal{R}_{G}$ $\supseteq\bigcup_{P_{V|U},P_{U|\tilde{\mathbf{X}}}}\Big{\{}(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:$
$R_{S}$ $\leq\min_{k}I(\mathbf{Y}_{k};U|V)-\max_{l}I(\mathbf{Z}_{l};U|V),$
$R_{J}$ $\geq\max_{k}I(\tilde{\mathbf{X}};U|V,\mathbf{Y}_{k})+\max_{k}I(\tilde{\mathbf{X}};V|\mathbf{Y}_{k}),$
$R_{L}$ $\geq\max_{k}I(\tilde{\mathbf{X}};U|V,\mathbf{Y}_{k})+\max_{k}I(\tilde{\mathbf{X}};V|\mathbf{Y}_{k})$
$~{}-I(\tilde{\mathbf{X}};U|X)+\min_{k}I(\mathbf{Y}_{k};V)-\min_{l}I(\mathbf{Z}_{l};V)\Big{\}},$ (7)
$\mathcal{R}_{C}$ $\supseteq\bigcup_{P_{V|U},P_{U|\tilde{\mathbf{X}}}}\Big{\{}(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:$
$R_{S}$ $\leq\min_{k}I(\mathbf{Y}_{k};U|V)-\max_{l}I(\mathbf{Z}_{l};U|V),$
$R_{J}$ $\geq\max_{k}I(\tilde{\mathbf{X}};U|V,\mathbf{Y}_{k})+\max_{k}I(\tilde{\mathbf{X}};V|\mathbf{Y}_{k})$
$~{}+\min_{k}I(\mathbf{Y}_{k};U|V)-\max_{l}I(\mathbf{Z}_{l};U|V),$
$R_{L}$ $\geq\max_{k}I(\tilde{\mathbf{X}};U|V,\mathbf{Y}_{k})+\max_{k}I(\tilde{\mathbf{X}};V|\mathbf{Y}_{k})$
$~{}-I(\tilde{\mathbf{X}};U|X)+\min_{k}I(\mathbf{Y}_{k};V)-\min_{l}I(\mathbf{Z}_{l};V)\Big{\}},$ (8)
where auxiliary random variables $V$ and $U$ satisfy the Markov chain $V-U-\tilde{\mathbf{X}}-X-(\mathbf{Y}_{k},\mathbf{Z}_{l})$ for all $k\in\mathcal{K}$ and $l\in\mathcal{L}$, and $|\mathcal{V}|\leq|\tilde{\mathbfcal{X}}|+6$ and $|\mathcal{U}|\leq(|\tilde{\mathbfcal{X}}|+6)(|\tilde{\mathbfcal{X}}|+3)$.
Proof.
See Appendix A-A. ∎
A special case of the GS model considered in this paper was investigated in [1]. One can check that, when $\Omega_{\tilde{X}}=\Omega_{Y}=\Omega_{Z}=|\mathcal{L}|=1$, the region in (7) reduces to [1, Prop. 1].
Proposition 2 (Outer bounds).
We have
$\mathcal{R}_{G}\subseteq$ $\bigcap_{k\in\mathcal{K}}\bigcap_{l\in\mathcal{L}}\bigcup_{P_{V|U},P_{U|\tilde{\mathbf{X}}}}\Big{\{}(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:~{}$
$R_{S}$ $\leq I(\mathbf{Y}_{k};U|V)-I(\mathbf{Z}_{l};U|V),$
$R_{J}$ $\geq I(\tilde{\mathbf{X}};U|\mathbf{Y}_{k}),$
$R_{L}$ $\geq I(X;U|\mathbf{Y}_{k})+I(\mathbf{Y}_{k};V)-I(\mathbf{Z}_{l};V)\Big{\}},$ (9)
$\mathcal{R}_{C}\subseteq$ $\bigcap_{k\in\mathcal{K}}\bigcap_{l\in\mathcal{L}}\bigcup_{P_{V|U},P_{U|\tilde{\mathbf{X}}{}}}\Big{\{}(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:~{}$
$R_{S}$ $\leq I(\mathbf{Y}_{k};U|V)-I(\mathbf{Z}_{l};U|V),$
$R_{J}$ $\geq I(\tilde{\mathbf{X}};U|\mathbf{Y}_{k})+I(\mathbf{Y}_{k};U|V)-I(\mathbf{Z}_{l};U|V),$
$R_{L}$ $\geq I(X;U|\mathbf{Y}_{k})+I(\mathbf{Y}_{k};V)-I(\mathbf{Z}_{l};V)\Big{\}},$ (10)
where $V$ and $U$ satisfy the Markov chain $V-U-\tilde{\mathbf{X}}-X-(\mathbf{Y}_{k},\mathbf{Z}_{l})$, and $|\mathcal{V}|\leq|\tilde{\mathbfcal{X}}|+5$, $|\mathcal{U}|\leq(|\tilde{\mathbfcal{X}}|+5)(|\tilde{\mathbfcal{X}}|+3)$.
Proof.
See Appendix A-B. ∎
The region in (9) matches [1, Prop. 2] when $\Omega_{\tilde{X}}=\Omega_{Y}=\Omega_{Z}=|\mathcal{L}|=1$. Moreover, in the non-compound settings, i.e., when $|\mathcal{K}|=1=|\mathcal{L}|$, the bounds in Propositions 1 and 2 match, yielding the following result.
Corollary 1 (Capacity regions).
When $|\mathcal{K}|=1=|\mathcal{L}|$, we have
$\mathcal{R}_{G}=$ $\bigcup_{P_{V|U},P_{U|\tilde{\mathbf{X}}}}\Big{\{}(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:~{}$
$R_{S}$ $\leq I(\mathbf{Y};U|V)-I(\mathbf{Z};U|V),$
$R_{J}$ $\geq I(\tilde{\mathbf{X}};U|\mathbf{Y}),$
$R_{L}$ $\geq I(X;U|\mathbf{Y})+I(\mathbf{Y};V)-I(\mathbf{Z};V)\Big{\}},$ (11)
$\mathcal{R}_{C}=$ $\bigcup_{P_{V|U},P_{U|\tilde{\mathbf{X}}{}}}\Big{\{}(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:~{}$
$R_{S}$ $\leq I(\mathbf{Y};U|V)-I(\mathbf{Z};U|V),$
$R_{J}$ $\geq I(\tilde{\mathbf{X}};U|\mathbf{Y})+I(\mathbf{Y};U|V)-I(\mathbf{Z};U|V),$
$R_{L}$ $\geq I(X;U|\mathbf{Y})+I(\mathbf{Y};V)-I(\mathbf{Z};V)\Big{\}}.$ (12)
where $(U,V)$ satisfy the same conditions as in Proposition 2.
Proof.
We only sketch the proof of (11), as that of (12) follows similarly. When $|\mathcal{K}|=|\mathcal{L}|=1$ and by dropping the indices $k$ and $l$, the constraints in Proposition 1 become
$R_{S}$ $\leq I(\mathbf{Y};U|V)-I(\mathbf{Z};U|V),$ (13)
$R_{J}$ $\geq I(\tilde{\mathbf{X}};U|V,\mathbf{Y})+I(\tilde{\mathbf{X}};V|\mathbf{Y})\overset{\rm(a)}{=}I(\tilde{\mathbf{X}};U|\mathbf{Y}),$ (14)
$R_{L}$ $\geq I(\tilde{\mathbf{X}};U|V,\mathbf{Y})+I(\tilde{\mathbf{X}};V|\mathbf{Y})$
$~{}~{}~{}-I(\tilde{\mathbf{X}};U|X)+I(\mathbf{Y};V)-I(\mathbf{Z};V)$
$=I(\tilde{\mathbf{X}};U|\mathbf{Y})-I(\tilde{\mathbf{X}};U|X)+I(\mathbf{Y};V)-I(\mathbf{Z};V)$
=$\overset{\rm(b)}{=}I(X;U|\mathbf{Y})+I(\mathbf{Y};V)-I(\mathbf{Z};V),$ (15)
where (a) and (b) hold by the Markov chains $V-U-\tilde{\mathbf{X}}-\mathbf{Y}$ and $U-\tilde{\mathbf{X}}-X-\mathbf{Y}$, respectively. As $|\mathcal{K}|=|\mathcal{L}|=1$, (13)–(15) match (9) in Proposition 2, and thus (11) is proved. ∎
Remark 1.
The privacy-leakage rate in [20, Th. 3 and 4] without action cost is bounded as
$R_{L}$ $\geq I(X;U,Y)-I(X;Y|V)+I(X;Z|V)$
$=I(X;U|Y)+I(Y;V)-I(Z;V)+I(X;Z),$
where the equality holds by the Markov chain $V-X-(Y,Z)$. Compared to the privacy-leakage rate in (11) and (12), there is an extra term $I(X;Z)$, because the privacy-leakage rate constraint in [20, Th. 3 and 4] is defined as $\frac{1}{n}I(X^{n};J,Z^{n})=\frac{1}{n}I(X^{n};J|Z^{n})+I(X;Z)$. Therefore, (11) and (12) coincide with [20, Th. 3 and 4] (without action cost) if [20, eq. (5)] is replaced by (6).
In general, the regions in Propositions 1 and 2 do not coincide. More specifically, the inner bounds are obtained by optimizing the test channels $P_{V|U}$ and $P_{U|\tilde{\mathbf{X}}}$ for all possible channel states, while for the outer bounds, the intersection over the indices $k$ and $l$ is taken before the union of the test channels. Optimizing for an arbitrary state index pair could yield a larger region compared to the one for all possible channel pairs, leading to a gap between the inner and outer bounds, similar to the observation drawn for compound-wiretap channels in [34].
In the next subsection, we demonstrate that the regions in Propositions 1 and 2 match for Gaussian sources.
III-B Gaussian Sources
In this subsection, we limit our discussion to a special case of setup in Section III-A where the enrollment channel is noiseless, i.e., $\tilde{X}=X$. We consider $P_{X\mathbf{Y}_{k}\mathbf{Z}_{l}}$ the joint distribution of zero-mean Gaussian random variables with a non-singular covariance matrix. Suppose that the source $X\sim\mathcal{N}(0,\sigma^{2}_{X})$, then it suffices to model the channels to the decoder and Eve as follows.
Lemma 1.
Without loss of generality, one can write
$\mathbf{Y}_{k}=\mathbf{H}_{k}X+\mathbf{N}_{\mathbf{Y}_{k}},~{}\mathbf{Z}_{l}=\mathbf{H}_{l}X+\mathbf{N}_{\mathbf{Z}_{l}},$ (16)
where $\mathbf{H}_{k}\in\mathbb{R}^{\Omega_{Y}\times 1}$, $\tilde{\mathbf{H}}_{l}\in\mathbb{R}^{\Omega_{Z}\times 1}$, and $\mathbf{N}_{\mathbf{Y}_{k}}\sim{\bf\mathcal{N}}({\bf 0},\mathbf{I}_{\Omega_{Y}})$, and $\mathbf{N}_{\mathbf{Z}_{l}}\sim{\bf\mathcal{N}}({\bf 0},\mathbf{I}_{\Omega_{Z}})$ are independent of $X$. Here, $\mathbf{I}$ denotes the identity matrix.
Proof.
See Appendix B. ∎
Note that the single-letter expressions characterized in Propositions 1 and 2 can be extended to the channel model (16). In proving the capacity results for the GS and CS models with Gaussian sources, we leverage Proposition 1 to show the achievability part. However, for the outer bounds, we transform the problem in (16) into a scalar Gaussian problem using sufficient statistics [39, Ch. 2]. After the transformation, it is unclear whether all constraints in Definition 3, particularly (2), remain preserved under the scalar random variables. As a result, Proposition 2 may not be directly applicable. To this end, as shown in the proof of Theorem 1, we derive new outer regions for the channel model (16) to establish the converse part of Theorem 1.
In the sequel, we define
$k^{*}\in\mathop{\rm argmin}\limits_{k\in\mathcal{K}}\{\mathbf{H}^{\intercal}_{k}\mathbf{H}_{k}\},~{}~{}~{}l^{*}\in\mathop{\rm argmax}\limits_{l\in\mathcal{L}}\{\tilde{\mathbf{H}}^{\intercal}_{l}\tilde{\mathbf{H}}_{l}\}.$ (17)
To simplify the presentation of the results for Gaussian sources, we define the following rate constraints:
$R_{S}$ $\leq\frac{1}{2}\log\left(\frac{(\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}+1)(\alpha\sigma^{2}_{X}\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\tilde{\mathbf{H}}_{l^{*}}+1)}{(\alpha\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}+1)(\sigma^{2}_{X}\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\tilde{\mathbf{H}}_{l^{*}}+1)}\right),$ (18)
$R_{J}$ $\geq\frac{1}{2}\log\left(\frac{\alpha\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}+1}{\alpha(\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}+1)}\right),$ (19)
$R_{J}$ $\geq\frac{1}{2}\log\left(\frac{\alpha\sigma^{2}_{X}\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\tilde{\mathbf{H}}_{l^{*}}+1}{\alpha(\sigma^{2}_{X}\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\tilde{\mathbf{H}}_{l^{*}}+1)}\right),$ (20)
$R_{L}$ $\geq\frac{1}{2}\log\left(\frac{\alpha\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}+1}{\alpha(\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}+1)}\right).$ (21)
Theorem 1 (Capacity regions).
If $\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}\geq\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\tilde{\mathbf{H}}_{l^{*}}$, then the capacity regions of the GS and CS models are
[…]
$~{}{\rm are~{}satisfied}\},$ (22)
[…]
$~{}{\rm are~{}satisfied}\}.$ (23)
If $\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}<\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\tilde{\mathbf{H}}_{l^{*}}$, then $\mathcal{R}_{G}=\mathcal{R}_{C}=\{(R_{S},R_{J},R_{L}):R_{S}=0,R_{J}\geq 0,R_{L}\geq 0\}.$
Proof.
See Appendix C. ∎
As a special case, when $\Omega_{Y}$, $\Omega_{Z}$, $|\mathcal{K}|$, and $|\mathcal{L}|$ are all one (let $\mathbf{H}=h$ and $\tilde{\mathbf{H}}=\tilde{h}$), the AWGN channels to the decoder and Eve reduce to $Y=hX+N_{Y}$ and $Z=\tilde{h}X+N_{Z}$, respectively, with $N_{Y}\sim\mathcal{N}(0,1)$ and $N_{Z}\sim\mathcal{N}(0,1)$. In this case, using the correlation coefficients of $(X,Y)$, $\rho^{2}_{XY}=\sigma^{2}_{X}h^{2}/(\sigma^{2}_{X}h^{2}+1)$ and that of $(X,Z)$, $\rho^{2}_{XZ}=\sigma^{2}_{X}\tilde{h}^{2}/(\sigma^{2}_{X}\tilde{h}^{2}+1)$, the regions (22) and (23) can be transformed as
$\mathcal{R}_{G}$ $:\displaystyle=\bigcup_{0<\alpha\leq 1}\Big{\{}(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}$ :
$R_{S}\leq\frac{1}{2}\log\frac{\alpha\rho^{2}_{XZ}+1-\rho^{2}_{XZ}}{\alpha\rho^{2}_{XY}+1-\rho^{2}_{XY}},$
$R_{J}\geq\frac{1}{2}\log\frac{\alpha\rho^{2}_{XY}+1-\rho^{2}_{XY}}{\alpha},$
$R_{L}\geq\frac{1}{2}\log\frac{\alpha\rho^{2}_{XY}+1-\rho^{2}_{XY}}{\alpha}\Big{\}},$ (24)
$\mathcal{R}_{C}$ $:\displaystyle=\bigcup_{0<\alpha\leq 1}\Big{\{}(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}$ :
$R_{S}\leq\frac{1}{2}\log\frac{\alpha\rho^{2}_{XZ}+1-\rho^{2}_{XZ}}{\alpha\rho^{2}_{XY}+1-\rho^{2}_{XY}},$
$R_{J}\geq\frac{1}{2}\log\frac{\alpha\rho^{2}_{XZ}+1-\rho^{2}_{XZ}}{\alpha},$
$R_{L}\geq\frac{1}{2}\log\frac{\alpha\rho^{2}_{XY}+1-\rho^{2}_{XY}}{\alpha}\Big{\}}.$ (25)
The regions in (24) and (25) align with [21, Cor. 1] when [21, eq. (5)] is replaced with (6) to eliminate the quantity $I(X;Z)$. Moreover, when the storage rate is not considered, i.e., (4) is not imposed, and Eve has no side information, i.e., $\rho_{XZ}=0$), the regions in (24) and (25) simplify to [10, Th. 4.1 and 4.2].
III-C Numerical Examples
We now present numerical examples for the capacity region of the GS model. Specifically, we consider three cases, with the parameters summarized as follows: (a) $\Omega_{Y}=\Omega_{Z}=1$ with $\mathbf{H}^{\intercal}_{k^{*}}=0.95$ and $\mathbf{H}^{\intercal}_{l^{*}}=0.8$, (b) $\Omega_{Y}=3$ and $\Omega_{Z}=1$ with $\mathbf{H}^{\intercal}_{k^{*}}=[0.95~{}0.95~{}0.95]$ and $\mathbf{H}^{\intercal}_{l^{*}}=0.8$, (c) $\Omega_{Y}=3$ and $\Omega_{Z}=4$ with $\mathbf{H}^{\intercal}_{k^{*}}=[0.95~{}0.95~{}0.95]$ and $\mathbf{H}^{\intercal}_{l^{*}}=[0.8~{}0.8~{}0.5~{}0.5]$. Moreover, we fix the variance of the source identifier as $\sigma^{2}_{X}=5$ for all cases.
For a given $\alpha$, define the optimal storage rate $R_{J}(\alpha)=\frac{1}{2}\log\frac{\alpha\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}+1}{\alpha(\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}+1)},$ from which one can express $\alpha$ as
$\alpha=\frac{1}{2^{2R_{J}(\alpha)}+(2^{2R_{J}(\alpha)}-1)\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}}.$ (26)
Substituting (26) into the right-hand side of (18), the optimal secret-key rate based on $R_{J}(\alpha)$ is given by $R_{S}(R_{J}(\alpha))=\frac{1}{2}\log\left(\frac{\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}(1-2^{-2{R_{J}(\alpha)}})+\sigma^{2}_{X}\mathbf{H}^{\intercal}_{l^{*}}\mathbf{H}_{l^{*}}2^{-2{R_{J}(\alpha)}}+1}{\sigma^{2}_{X}\mathbf{H}^{\intercal}_{l^{*}}\mathbf{H}_{l^{*}}+1}\right)$. Note that if $R_{J}(\alpha)\rightarrow\infty$, $R^{*}_{S}(R_{J}(\alpha))\rightarrow\frac{1}{2}\log\left(\frac{\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}+1}{\sigma^{2}_{X}\mathbf{H}^{\intercal}_{l^{*}}\mathbf{H}_{l^{*}}+1}\right)$.
Figure 3 depicts the relationship of $(R_{J}(\alpha),R_{S}({R_{J}(\alpha)}))$. In this figure, Case 2 (blue curve) shows a high secret-key rate compared to the other cases. This is due to an increase in the number of antennas at the decoder, which enhances the correlation between the source and observations at the terminal. On the other hand, in Case 3 (red curve), as the number of antennas at Eve increases, the secret-key rate drops compared to Case 2 because the stronger correlation with Eve reduces the key-generation rate. Also, this case shows that even when Eve has more antennas, a positive secret-key rate is still achievable as long as $\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}\geq\mathbf{H}^{\intercal}_{l^{*}}\mathbf{H}_{l^{*}}$.
IV Concluding Remarks
We studied the GS and CS models for secret-key generation from private identifiers when the authentication channels to the legitimate terminals and Eve have multiple states and outputs. In addition, the channel uncertainty for the decoder and Eve was also taken into consideration. We first derived inner and outer bounds for general discrete sources. Next, we focused on Gaussian sources, and gave full characterizations of the capacity regions in this case. Furthermore, numerical evaluations for the Gaussian case revealed how the number of antennas at the legitimate terminals and eavesdropper affects the key and storage rates.
Appendix A Proof of Propositions 1 and 2
A-A Proof of Proposition 1
We only prove (7) since (8) follows similarly with an extra procedure, a one-time pad procedure to conceal the chosen secret key. As a result, an extra rate equal to the secret-key rate is needed for storing the concealed key information in the database, which appears in the constraint of the storage rate of the CS model.
Fix the test channels $P_{U|\tilde{\mathbf{X}}}$ and $P_{V|U}$ and let $\delta>0$. In the following, we show that these rates are achievable
$R_{S}$ $~{}{\triangleq}~{}\min_{k}I(\mathbf{Y}_{k};U|V)-\max_{l}I(\mathbf{Z}_{l};U|V)-\delta,$
$R_{J}$ $~{}{\triangleq}~{}\max_{k}I(\tilde{\mathbf{X}};U|V,\mathbf{Y}_{k})+\max_{k}I(\tilde{\mathbf{X}};V|\mathbf{Y}_{k})+5\delta,$
$R_{L}$ $~{}{\triangleq}~{}\max_{k}I(\tilde{\mathbf{X}};U|V,\mathbf{Y}_{k})+\max_{k}I(\tilde{\mathbf{X}};V|\mathbf{Y}_{k})$
$~{}~{}~{}-I(\tilde{\mathbf{X}};U|X)+\min_{k}I(\mathbf{Y}_{k};V)-\min_{l}I(\mathbf{Z}_{l};V)+4\delta.$
For the random codebook construction, we also define
$R_{v}$ ${\triangleq}~{}I(\tilde{\mathbf{X}};V)+\delta,~{}~{}{R_{J_{v_{1}}}}{\triangleq}\max_{k}I(\tilde{\mathbf{X}};V|\mathbf{Y}_{k})+2\delta,$
$R_{u}$ ${\triangleq}~{}I(\tilde{\mathbf{X}};U|V)+\delta,~{}~{}{R_{J_{u_{1}}}}{\triangleq}\max_{k}I(\tilde{\mathbf{X}};U|\mathbf{Y}_{k},V)+3\delta,$
and the sets $\mathcal{J}_{v_{1}}{\triangleq}[1:2^{n{R_{J_{v_{1}}}}}]$, $\mathcal{J}_{v_{2}}{\triangleq}[1:2^{n(R_{v}-{R_{J_{v_{1}}}})}]$, $\mathcal{J}_{u_{1}}{\triangleq}[1:2^{n{R_{J_{u_{1}}}}}]$, $\mathcal{J}_{u_{2}}{\triangleq}[1:2^{nR_{S}}]$, $\mathcal{J}_{u_{3}}{\triangleq}[1:2^{n(\max_{l}I(\mathbf{Z}_{l};U|V)-\delta)}]$. Note that $R_{u}={R_{J_{u_{1}}}}+R_{S}+\max_{l}I(\mathbf{Z}_{l};U|V)-\delta$.
Figure 3: The relation of storage and secret-key rates.
Random Codebook: Generate i.i.d. sequences $v^{n}(j_{v_{1}},j_{v_{2}})$ from $P_{V^{n}}$, where $(j_{v_{1}},j_{v_{2}})\in\mathcal{J}_{v_{1}}\times\mathcal{J}_{v_{2}}$. For every $(j_{v_{1}},j_{v_{2}})$, generate i.i.d. sequences $u^{n}(j_{u_{1}},{j_{u_{2}},j_{u_{3}}},j_{v_{1}},j_{v_{2}})$, where $(j_{u_{1}},j_{u_{2}},j_{u_{3}})\in\mathcal{J}_{u_{1}}\times\mathcal{J}_{u_{2}}\times\mathcal{J}_{u_{3}}$, according to $P_{U^{n}|V^{n}=v^{n}(j_{v_{1}},j_{v_{2}})}$. All the generated sequences $(V^{n}(j_{v_{1}},j_{v_{2}}),U^{n}(j_{u_{1}},j_{u_{2}},j_{u_{3}},j_{v_{1}},j_{v_{2}}))$ form the  codebook $\mathcal{C}_{n}$.
Encoding: Observing $\tilde{\mathbf{x}}^{n}{}$, the encoder first finds $(j_{v_{1}},j_{v_{2}})$ such that $(\tilde{\mathbf{x}}^{n},v^{n}(j_{v_{1}},j_{v_{2}}))\in\mathcal{T}^{n}_{\delta}$. Then, it looks for $(j_{u_{1}},j_{u_{2}},j_{u_{3}})$ such that $(\tilde{\mathbf{x}}^{n},u^{n}(j_{u_{1}},j_{u_{2}},j_{u_{3}},j_{v_{1}},j_{v_{2}}))\in\mathcal{T}^{n}_{\delta}(\tilde{\mathbf{X}}U|v^{n}(j_{v_{1}},j_{v_{2}}))$. If a unique tuple $(j_{u_{1}},j_{u_{2}},j_{u_{3}},j_{v_{1}},j_{v_{2}})$ is found, the encoder assigns the helper data $j=(j_{v_{1}},j_{u_{1}})$ and the secret key $s=j_{u_{2}}$. If multiple such tuples are found, the encoder selects one tuple uniformly at random and assigns $j=(j_{v_{1}},j_{u_{1}})$ and $s=j_{u_{2}}$. In case no such tuple exists, the encoder sets all $j_{v_{1}}$, $j_{v_{2}}$, $j_{u_{1}}$, $j_{u_{2}}$, and $j_{u_{3}}$ to be one and assigns $j=(1,1)$ and $s=1$.
Decoding: From $\mathbf{y}^{n}_{k}$ and $(j_{u_{1}},j_{v_{1}})$, the decoder first looks for the unique index $\hat{j}_{v_{2}}$ such that $(\mathbf{y}^{n}_{k},v^{n}(j_{v_{1}},\hat{j}_{v_{2}}))\in\mathcal{T}^{n}_{\delta}$. Then, it looks for the unique pair $(\hat{j}_{u_{2}},\hat{j}_{u_{3}})$ such that $(\mathbf{y}^{n}_{k},u^{n}(j_{u_{1}},\hat{j}_{u_{2}},\hat{j}_{u_{3}},j_{v_{1}},\hat{j}_{v_{2}}))\in\mathcal{T}^{n}_{\delta}(\mathbf{Y}_{k}U|v^{n}(j_{v_{1}},\hat{j}_{v_{2}}))$. If the indices $\hat{j}_{u_{2}},\hat{j}_{u_{3}}$, and $\hat{j}_{v_{2}}$ are uniquely determined, then the decoder estimates $\hat{s}=\hat{j}_{u_{2}}$; otherwise, it sets $\hat{s}=1$ and declares an error.
In the following, we write $V^{n}(J_{v_{1}},J_{v_{2}})$ and $U^{n}(J_{u_{1}},J_{u_{2}},J_{u_{3}},J_{v_{1}},J_{v_{2}})$ as $V^{n}$ and $U^{n}$ for convenience.
Analysis of Error Probability:   Possible error events at the encoder are
  • $\mathcal{E}_{1}$
    : $\{(\tilde{\mathbf{X}}^{n}{},V^{n}(j_{v_{1}},j_{v_{2}}))~{}{\notin}~{}\mathcal{T}^{n}_{\delta},\forall(j_{v_{1}},j_{v_{2}})\in\mathcal{J}_{v_{1}}\times\mathcal{J}_{v_{2}}\}$,
  • $\mathcal{E}_{2}$
    : $\{(\tilde{\mathbf{X}}^{n},U^{n}(j_{u_{1}},j_{u_{2}},j_{u_{3}},J_{v_{1}},J_{v_{2}}))\notin\mathcal{T}^{n}_{\delta}(\tilde{\mathbf{X}}U|V^{n})$, $\forall(j_{u_{1}},j_{u_{2}},j_{u_{3}})\in\mathcal{J}_{u_{1}}\times\mathcal{J}_{u_{2}}\times\mathcal{J}_{u_{3}}\}$,
and those at the decoder are
  • $\mathcal{E}_{3}$
    : $\{(\mathbf{Y}^{n}_{k},U^{n},V^{n})\notin\mathcal{T}^{n}_{\delta}\}$,
  • $\mathcal{E}_{4}$
    : $\{\exists j^{\prime}_{v_{2}}\in\mathcal{J}_{v_{2}},j^{\prime}_{v_{2}}\neq J_{v_{2}}$ and $(\mathbf{Y}^{n}_{k},V^{n}(J_{v_{1}},j^{\prime}_{v_{2}}))\in\mathcal{T}^{n}_{\delta}\}$,
  • $\mathcal{E}_{5}$
    : $\{\exists(j^{\prime}_{u_{2}},j^{\prime}_{u_{3}})\in\mathcal{J}_{u_{2}}\times\mathcal{J}_{u_{3}},(j^{\prime}_{u_{2}},j^{\prime}_{u_{3}})\neq(J_{u_{2}},J_{u_{3}})$ and $(\mathbf{Y}^{n}_{k},U^{n}(J_{u_{1}},j^{\prime}_{u_{2}},j^{\prime}_{u_{3}},J_{v_{1}},J_{v_{2}}))\in\mathcal{T}^{n}_{\delta}(\mathbf{Y}_{k}U|V^{n})\}$.
Then, we have
$\max_{k}{\mathbb{P}}\{\hat{S}_{k}\neq S\}=\mathbb{P}\{\cup_{i=1}^{5}\mathcal{E}_{i}\}$
$\leq\mathbb{P}\{\mathcal{E}_{1}\}+\mathbb{P}\{\mathcal{E}_{2}\}+\mathbb{P}\{\mathcal{E}_{3}\cap(\mathcal{E}_{1}\cup\mathcal{E}_{2})^{c}\}+\mathbb{P}\{\mathcal{E}_{4}\}+\mathbb{P}\{\mathcal{E}_{5}\}.$
The first and second terms vanish by the covering lemma [43, Lemma 3.3] since $R_{v}>I(\tilde{\mathbf{X}};V)$ and $R_{u}>I(\tilde{\mathbf{X}};U|V)$, respectively. The third term vanishes by Markov lemma [39, Lemma 15.8.1]. The last two terms vanish by the packing lemma [43, Lemma 3.1], since the rate of index $\hat{j}_{v_{2}}$ is less than $\min_{k}I(\mathbf{Y}_{k};V)$ and that of index pair $(\hat{j}_{u_{2}},\hat{j}_{u_{3}})$ is less than $\min_{k}I(\mathbf{Y}_{k};U|V)$, respectively. Hence, we have
$\lim_{n\rightarrow\infty}\max_{k}{\mathbb{P}}\{\hat{S}_{k}\neq S\}\rightarrow 0.$ (27)
Before we analyze the constraints (3), (4), (5), and (6) in Definition 3, we state the following lemmas.
Lemma 2.
(Adapted from [34, Lemma A.1]) If the inequality $I(\mathbf{Z}_{l};U|V)<I(\mathbf{Z}_{l^{\prime}};U|V)$ holds for $l,l^{\prime}\in\mathcal{L}$, then there exists a random variable $A$ such that the equality $I(\mathbf{Z}_{l},A;U|V)=I(\mathbf{Z}_{l^{\prime}};U|V)$ and the Markov chain $V-U-\tilde{\mathbf{X}}-(\mathbf{Z}_{l},\mathbf{Z}_{l^{\prime}})-A$ are satisfied.
An Intermediate Step: For any $l\in\mathcal{L}$, by Lemma 2, there exists $A$ such that for
$\tilde{\mathbf{Z}}\triangleq(Z_{l},A),$ (28)
we have $I(\tilde{\mathbf{Z}};U|V)=\max_{l}I(\mathbf{Z}_{l};U|V)$ and
$V-U-\tilde{\mathbf{X}}-\tilde{\mathbf{Z}}.$ (29)
Moreover, define a binary random variable $T$, which takes $1$ if $(U^{n},\tilde{\mathbf{X}}^{n},\tilde{\mathbf{Z}}^{n})\in\mathcal{T}^{n}_{\delta}$ and 00 otherwise. For large enough $n$, it holds that
$P_{T}(1)\geq 1-\tilde{\delta}_{n}$ (30)
with $\tilde{\delta}_{n}\downarrow 0$ as $\delta\downarrow 0$ and $n\rightarrow\infty$. This follows because the pair $(U^{n},\tilde{\mathbf{X}}^{n})$ is jointly typical with probability approaching one, as shown in (27), and $\tilde{\mathbf{Z}}^{n}$ is i.i.d. generated according to $\prod_{t=1}^{n}P_{\tilde{\mathbf{Z}}_{t}|\tilde{\mathbf{X}}_{t}}$ from (29), and thus (30) follows by applying the Markov lemma [39, Lemma 15.8.1]. Similarly, we have joint typicality of $(V^{n},\tilde{\mathbf{Z}}^{n})$ as $(V^{n},\tilde{\mathbf{X}}^{n})$ is jointly typical with high probability. These properties are applied in proving the next lemma, which plays a key role in the analyses of the secret-key uniformity and secrecy-leakage.
Lemma 3.
For an arbitrary index $l\in\mathcal{L}$, we have
$H(J_{u_{2}}|J_{u_{1}},J_{v_{1}},\mathbf{Z}^{n}_{l},\mathcal{C}_{n})$ $\geq n(R_{S}-\xi_{n}),$ (31)
where $\xi_{n}$ goes to zero as $\delta\downarrow 0$ and $n\rightarrow\infty$.
Proof.
We have
$H(J_{u_{2}}|J_{u_{1}},J_{v_{1}},\mathbf{Z}^{n}_{l},\mathcal{C}_{n})$
$\geq H(J_{u_{2}}|J_{u_{1}},J_{v_{1}},J_{v_{2}},\mathbf{Z}^{n}_{l},A^{n},\mathcal{C}_{n})$
=$\overset{\rm(a)}{=}H(J_{u_{2}}|J_{u_{1}},J_{v_{1}},J_{v_{2}},\tilde{\mathbf{Z}}^{n},\mathcal{C}_{n})$
$=H(J_{u_{1}},J_{u_{2}},J_{u_{3}},J_{v_{1}},J_{v_{2}},\tilde{\mathbf{Z}}^{n}|\mathcal{C}_{n})$
$~{}~{}~{}-H(J_{u_{3}}|J_{u_{1}},J_{u_{2}},J_{v_{1}},J_{v_{2}},\tilde{\mathbf{Z}}^{n},\mathcal{C}_{n})$
$~{}~{}~{}-H(J_{u_{1}},J_{v_{1}},J_{v_{2}},\tilde{\mathbf{Z}}^{n}|\mathcal{C}_{n})$
$\overset{\rm(b)}{\geq}H(J_{u_{1}},J_{u_{2}},J_{u_{3}},J_{v_{1}},J_{v_{2}},\tilde{\mathbf{Z}}^{n}|\mathcal{C}_{n})$
$~{}~{}~{}-H(J_{u_{1}},J_{v_{1}},J_{v_{2}},\tilde{\mathbf{Z}}^{n}|\mathcal{C}_{n})-n\delta_{n}$
$\overset{\rm(c)}{\geq}H(U^{n},\tilde{\mathbf{Z}}^{n}|\mathcal{C}_{n})-H(\tilde{\mathbf{Z}}^{n}|V^{n},\mathcal{C}_{n})$
$~{}~{}~{}-H(J_{u_{1}},J_{v_{1}},J_{v_{2}}|\mathcal{C}_{n})-n\delta_{n}$
$\geq~{}P_{T}(1)H(U^{n},\tilde{\mathbf{Z}}^{n}|T=1,\mathcal{C}_{n})-H(\tilde{\mathbf{Z}}^{n}|V^{n},\mathcal{C}_{n})$
$~{}~{}~{}{-H(J_{u_{1}}|\mathcal{C}_{n})-H(J_{v_{1}}|\mathcal{C}_{n})-H(J_{v_{2}}|\mathcal{C}_{n})-n\delta_{n}}$
$\overset{\rm(d)}{\geq}~{}(1-\tilde{\delta}_{n})H(U^{n},\tilde{\mathbf{Z}}^{n}|T=1,\mathcal{C}_{n})-H(\tilde{\mathbf{Z}}^{n}|V^{n},\mathcal{C}_{n})$
$~{}~{}~{}{-H(J_{u_{1}}|\mathcal{C}_{n})-H(J_{v_{1}}|\mathcal{C}_{n})-H(J_{v_{2}}|\mathcal{C}_{n})-n\delta_{n}}$
$\overset{\rm(e)}{\geq}n(1-\tilde{\delta}_{n})(I(\tilde{\mathbf{X}};U)+H(\tilde{\mathbf{Z}}|U)-2\epsilon_{\delta})-H(\tilde{\mathbf{Z}}^{n}|V^{n},\mathcal{C}_{n})$
$~{}~{}~{}-H(J_{u_{1}}|\mathcal{C}_{n})-H(J_{v_{1}}|\mathcal{C}_{n})-H(J_{v_{2}}|\mathcal{C}_{n})-n\delta_{n}$
$\overset{\rm(f)}{\geq}n(I(\tilde{\mathbf{X}};U)+H(\tilde{\mathbf{Z}}|U)-H(\tilde{\mathbf{Z}}|V)-\gamma_{n})$
$~{}~{}~{}-H(J_{u_{1}}|\mathcal{C}_{n})-H(J_{v_{1}}|\mathcal{C}_{n})-H(J_{v_{2}}|\mathcal{C}_{n})-{n\delta^{\prime}_{n}}$
$\overset{\rm(g)}{\geq}n(I(\tilde{\mathbf{X}};U)-I(\tilde{\mathbf{Z}};U|V)-\gamma_{n})$
$~{}~{}~{}-n(\max_{k}I(\tilde{\mathbf{X}};U|V,\mathbf{Y}_{k})+3\delta)$
$~{}~{}~{}-n(\max_{k}I(\tilde{\mathbf{X}};V|\mathbf{Y}_{k})+2\delta)$
$~{}~{}~{}-n(\min_{k}I(\mathbf{Y}_{k};V)-\delta)-n\delta^{\prime}_{n}$
=$\overset{\rm(h)}{=}n(I(\tilde{\mathbf{X}};U)-I(\tilde{\mathbf{Z}};U|V))$
$~{}~{}~{}-n(\max_{k}\{I(\tilde{\mathbf{X}};U|V)-I(\mathbf{Y}_{k};U|V)\})$
$~{}~{}~{}-n(\max_{k}\{I(\tilde{\mathbf{X}};V)-I(\mathbf{Y}_{k};V)\})-n\min_{k}I(\mathbf{Y}_{k};V)$
$~{}~{}~{}-n(4\delta+\gamma_{n}+\delta^{\prime}_{n})$
=$\overset{\rm(i)}{=}n(\min_{k}I(\mathbf{Y}_{k};U|V)-\max_{l}I(\mathbf{Z}_{l};U|V)-\delta-\xi_{n})$
$=n(R_{S}-\xi_{n}),$ (32)
where (a) holds from (28), (b) follows because the index $J_{u_{3}}$ can be reliably estimated from $(J_{u_{1}},J_{u_{2}},J_{v_{1}},J_{v_{2}},\tilde{\mathbf{Z}}^{n})$, as $\frac{1}{n}\log|\mathcal{J}_{3}|<\max_{l}I(\mathbf{Z}_{l};U|V)=I(\tilde{\mathbf{Z}};U|V)$, (c) holds because $U^{n}$ and $V^{n}$ are determined by the tuple $(J_{u_{1}},J_{u_{2}},J_{u_{3}},J_{v_{1}},J_{v_{2}})$ and the pair $(J_{v_{1}},J_{v_{2}})$, respectively, (d) follows from (30), and (e) follows from
$P_{\tilde{\mathbf{Z}}^{n}U^{n}}(\tilde{\mathbf{z}}^{n},u^{n})$ $\leq\sum_{\tilde{\mathbf{x}}^{n}\in\mathcal{T}^{n}_{\delta}(\tilde{\mathbf{X}}|\tilde{\mathbf{z}}^{n},u^{n})}P_{\tilde{\mathbf{X}}^{n}\tilde{\mathbf{Z}}^{n}}(\tilde{\mathbf{x}}^{n},\tilde{\mathbf{z}}^{n})$
$\leq 2^{n(H(\tilde{\mathbf{X}}|\tilde{\mathbf{Z}},U)-\epsilon_{\delta})}\cdot 2^{-n(H(\tilde{\mathbf{X}},\tilde{\mathbf{Z}})-\epsilon_{\delta})}$
$=2^{-n(I(\tilde{\mathbf{X}};U)+H(\tilde{\mathbf{Z}}|U)-2\epsilon_{\delta})},$
(f) follows because, as shown in the intermediate step, $(V^{n},\tilde{\mathbf{Z}}^{n})$ is jointly typical with high probability and thus $H(\tilde{\mathbf{Z}}^{n}|V^{n},\mathcal{C}_{n})\leq n(H(\tilde{\mathbf{Z}}|V)+\gamma_{n})$ (cf. [44, eq. (16)]) and $\delta^{\prime}_{n}\triangleq\tilde{\delta}_{n}(I(\tilde{\mathbf{X}};U)+H(\tilde{\mathbf{Z}}|U))+2(1-\tilde{\delta}_{n})\epsilon_{\delta}+\delta_{n}$, (g) is due to the Markov chain (29) and $H(J_{u_{1}}|\mathcal{C}_{n})\leq n{R_{J_{u_{1}}}}$, $H(J_{v_{1}}|\mathcal{C}_{n})\leq~{}n{R_{J_{v_{1}}}}$, $H(J_{v_{2}}|\mathcal{C}_{n})\leq n(R_{v}-R_{J_{v_{1}}})=n(\min_{k}I(\mathbf{Y}_{k};V)-\delta)$, (h) is due to the Markov chain $V-U-\tilde{\mathbf{X}}-\mathbf{Y}_{k}$, (i) follows from $I(\tilde{\mathbf{Z}};U|V)=\max_{l}I(\mathbf{Z}_{l};U|V)$ and $\xi_{n}~{}\triangleq~{}3\delta+\gamma_{n}+{\delta^{\prime}_{n}}$. ∎
Analyses of Uniformity and Secrecy-leakage:   The constraints of (3) and (5) can be evaluated as
$H(S|\mathcal{C}_{n})$ $=H(J_{u_{2}}|\mathcal{C}_{n})$
$\geq H(J_{u_{2}}|J_{u_{1}},J_{v_{1}},\mathbf{Z}^{n}_{l},\mathcal{C}_{n})$
$\geq n(R_{S}-\xi_{n}),$ (33)
and
$\max_{l}$ $I(S;J,\mathbf{Z}^{n}_{l}|\mathcal{C}_{n})=\max_{l}I(J_{u_{2}};J_{u_{1}},J_{v_{1}},\mathbf{Z}^{n}_{l}|\mathcal{C}_{n})$
$=\max_{l}\{H(J_{u_{2}}|\mathcal{C}_{n})-H(J_{u_{2}}|J_{u_{1}},J_{v_{1}},\mathbf{Z}^{n}_{l},\mathcal{C}_{n})\}$
$\leq\max_{l}\{nR_{S}-n(R_{S}-\xi_{n})\}=n\xi_{n},$ (34)
where (33) and (34) follow from Lemma 3.
Analysis of Storage Rate:   The helper data is $J=(J_{v_{1}},J_{u_{1}})$, and thus the total storage rate is $\frac{1}{n}\log|\mathcal{J}_{v_{1}}||\mathcal{J}_{u_{1}}|=R_{J_{v_{1}}}+R_{J_{u_{1}}}=R_{J}.$
Analysis of Privacy-Leakage Rate: We have
$\max_{l}I(X^{n};J|\mathbf{Z}^{n}_{l},\mathcal{C}_{n})=\max_{l}I(X^{n};J_{u_{1}},J_{v_{1}}|\mathbf{Z}^{n}_{l},\mathcal{C}_{n})$
$=I(X^{n};J_{u_{1}},J_{v_{1}}|\mathcal{C}_{n})-\min_{l}I(\mathbf{Z}^{n}_{l};J_{u_{1}},J_{v_{1}}|\mathcal{C}_{n})$
$\leq H(J_{u_{1}}|\mathcal{C}_{n})+H(J_{v_{1}}|\mathcal{C}_{n})-H(J_{u_{1}},J_{v_{1}}|X^{n},\mathcal{C}_{n})$
$~{}~{}~{}-\min_{l}I(\mathbf{Z}^{n}_{l};J_{u_{1}},J_{v_{1}}|\mathcal{C}_{n})$
$=H(J_{u_{1}}|\mathcal{C}_{n})+H(J_{v_{1}}|\mathcal{C}_{n})-H(\tilde{\mathbf{X}}^{n},J_{u_{1}},J_{v_{1}}|X^{n},\mathcal{C}_{n})$
$~{}~{}~{}+H(\tilde{\mathbf{X}}^{n}|X^{n},J_{u_{1}},J_{v_{1}},\mathcal{C}_{n})-\min_{l}I(\mathbf{Z}^{n}_{l};J_{u_{1}},J_{v_{1}}|\mathcal{C}_{n})$
$\overset{\rm(a)}{\leq}H(J_{u_{1}}|\mathcal{C}_{n})+H(J_{v_{1}}|\mathcal{C}_{n})-nH(\tilde{\mathbf{X}}|X)$
$~{}~{}~{}+n(H(\tilde{\mathbf{X}}|X,U)+\epsilon^{\prime}_{n})-\min_{l}I(\mathbf{Z}^{n}_{l};J_{u_{1}},J_{v_{1}}|\mathcal{C}_{n})$
$\leq H(J_{u_{1}}|\mathcal{C}_{n})+H(J_{v_{1}}|\mathcal{C}_{n})-nI(\tilde{\mathbf{X}};U|X)$
$~{}~{}~{}-\min_{l}I(\mathbf{Z}^{n}_{l};J_{v_{1}}|\mathcal{C}_{n})+n\epsilon^{\prime}_{n}$
$\leq H(J_{u_{1}}|\mathcal{C}_{n})+H(J_{v_{1}}|\mathcal{C}_{n})-nI(\tilde{\mathbf{X}};U|X)$
$~{}~{}~{}-{\min_{l}\{H(\mathbf{Z}^{n}_{l})-H(\mathbf{Z}^{n}_{l}|J_{v_{1}},J_{v_{2}},\mathcal{C}_{n})\}}$
$~{}~{}~{}+\max_{l}I(J_{v_{2}};\mathbf{Z}^{n}_{l}|J_{v_{1}},\mathcal{C}_{n})+n\epsilon^{\prime}_{n}$
$\overset{\rm(b)}{\leq}H(J_{u_{1}}|\mathcal{C}_{n})+H(J_{v_{1}}|\mathcal{C}_{n})-nI(\tilde{\mathbf{X}};U|X)$
$~{}~{}~{}-{\min_{l}\{H(\mathbf{Z}^{n}_{l})-H(\mathbf{Z}^{n}_{l}|V^{n},\mathcal{C}_{n})\}}+H(J_{v_{2}}|\mathcal{C}_{n})+n\epsilon^{\prime}_{n}$
$\overset{\rm(c)}{\leq}H(J_{u_{1}}|\mathcal{C}_{n})+H(J_{v_{1}}|\mathcal{C}_{n})-nI(\tilde{\mathbf{X}};U|X)$
$~{}~{}~{}-n(\min_{l}\{H(\mathbf{Z}_{l})-H(\mathbf{Z}_{l}|V)\})+H(J_{v_{2}}|\mathcal{C}_{n})+n\epsilon^{\prime\prime}_{n}$
$\leq n(\max_{k}I(\tilde{\mathbf{X}};V|\mathbf{Y}_{k})+\max_{k}I(\tilde{\mathbf{X}};U|V,\mathbf{Y}_{k})+4\delta$
$~{}~{}~{}-I(\tilde{\mathbf{X}};U|X)+\min_{k}I(\mathbf{Y}_{k};V)-\min_{l}I(\mathbf{Z}_{l};V)+\epsilon^{\prime\prime}_{n})$
$=n(R_{L}+\epsilon^{\prime\prime}_{n}),$ (35)
where (a) follows from (36), shown below, and the codebook $\mathcal{C}_{n}$ is independent of $(\tilde{\mathbf{X}}^{n},X^{n},\mathbf{Y}^{n}_{k},\mathbf{Z}^{n}_{l})$, (b) holds because $V^{n}$ is a function of $(J_{v_{1}},J_{v_{2}})$, and (c) follows because $H(Z^{n}_{l}|V^{n},\mathcal{C}_{n})\leq n(H(Z_{l}|V)+\gamma^{\prime}_{n})$ and $\epsilon^{\prime\prime}_{n}~{}\triangleq~{}\gamma^{\prime}_{n}+\epsilon^{\prime}_{n}$. For brevity, define $E~{}\triangleq~{}(J_{u_{2}},J_{u_{3}},J_{v_{2}})$, where the decoder can reliably estimate the index $E$ for given $(\mathbf{Y}^{n}_{k},J_{u_{1}},J_{v_{1}})$. Observe that
$H(\tilde{\mathbf{X}}^{n}|X^{n},J_{u_{1}},J_{v_{1}},\mathcal{C}_{n})$
$=H(\tilde{\mathbf{X}}^{n}|X^{n},J_{u_{1}},J_{v_{1}},E,\mathcal{C}_{n})+I(E;\tilde{\mathbf{X}}^{n}|X^{n},J_{u_{1}},J_{v_{1}},{\mathcal{C}_{n}})$
$\overset{\rm(a)}{\leq}H(\tilde{\mathbf{X}}^{n}|X^{n},U^{n},\mathcal{C}_{n})+H(E|X^{n},J_{u_{1}},J_{v_{1}},{\mathcal{C}_{n}})$
$\overset{\rm(b)}{\leq}H(\tilde{\mathbf{X}}^{n}|X^{n},U^{n},\mathcal{C}_{n})+H(E|\mathbf{Y}^{n}_{k},J_{u_{1}},J_{v_{1}},{\mathcal{C}_{n}})$
$\overset{\rm(c)}{\leq}H(\tilde{\mathbf{X}}^{n}|X^{n},U^{n},\mathcal{C}_{n})+n\epsilon_{n}$
$\overset{\rm(d)}{\leq}n(H(\tilde{\mathbf{X}}|X,U)+\epsilon^{\prime}_{n}),$ (36)
where (a) follows since $U^{n}$ is a function of $(J_{u_{1}},J_{u_{2}},J_{u_{3}},J_{v_{1}},J_{v_{2}})$, (b) is due to the Markov chain $E-(X^{n},J_{u_{1}},J_{v_{1}})-\mathbf{Y}^{n}_{k}$ and conditioning reduces entropy, (c) follows from Fano’s inequality with $\epsilon_{n}\downarrow 0$ as $\delta\downarrow 0$ and $n\rightarrow\infty$, and (d) because $H(\tilde{\mathbf{X}}^{n}|X^{n},U^{n},\mathcal{C}_{n})\leq n(H(\tilde{\mathbf{X}}|X,U)+\gamma^{\prime\prime}_{n})$ for jointly typical sequences (cf. [44, eq. (16)]) and $\epsilon^{\prime}_{n}~{}{\triangleq}~{}\gamma^{\prime\prime}_{n}+\epsilon_{n}$.
From (27), (33), (34), and (35), there must be at least one codebook satisfying all the conditions in Definition 3, so that the region in (7) is achievable.
A-B Proof of Proposition 2
The cardinality bounds of the auxiliary random variables can be obtained from the support lemma [43, Lemma 3.4].
We show the proof for the GS model via a result derived in [20] in the absence of action cost. By replacing (6) with $I(X^{n};J,\mathbf{Z}^{n}_{l})\leq n(R_{L}+\delta)$, the outer bound for a pair $(k,l)$, denoted by $\mathcal{O}_{G_{kl}}$, is [20, Th. 3]
$\mathcal{O}_{G_{kl}}$ $~{}{\triangleq}\bigcup_{P_{V|U},P_{U|\tilde{\mathbf{X}}}}\{(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:~{}$
$R_{S}$ $\leq I(\mathbf{Y}_{k};U|V)-I(\mathbf{Z}_{l};U|V),~{}R_{J}\geq I(\tilde{\mathbf{X}};U|{\mathbf{Y}_{k}}),$
$R_{L}$ $\geq I(X;U,\mathbf{Y}_{k})-I(X;\mathbf{Y}_{k}|V)+I(X;\mathbf{Z}_{l}|V)\},$ (37)
where $V$ and $U$ satisfy $V-U-\tilde{\mathbf{X}}-X-(\mathbf{Y}_{k},\mathbf{Z}_{l})$.
In Definition 3, the constraints on the secret-key and storage rates are the same as in [20, Def. 6], and the resulting bounds are given in the same form as in (37). For the privacy-leakage rate, we expand the right-hand side of $R_{L}$ in (37) as
$I(X;U,\mathbf{Y}_{k})-I(X;\mathbf{Y}_{k}|V)+I(X;\mathbf{Z}_{l}|V)$
$=I(X;U|\mathbf{Y}_{k})+I(\mathbf{Y}_{k};V)-I(\mathbf{Z}_{l};V)+I(X;\mathbf{Z}_{l}),$ (38)
where we used the Markov chain $V-X-(\mathbf{Y}_{k},\mathbf{Z}_{l})$. By (38), the privacy-leakage rate is lower bounded as
$n(R_{L}+\delta)\geq I(X^{n};J|\mathbf{Z}^{n}_{l})=I(X^{n};J,\mathbf{Z}^{n}_{l})-I(X^{n};\mathbf{Z}^{n}_{l})$
$\geq n(I(X;U|\mathbf{Y}_{k})+I(\mathbf{Y}_{k};V)-I(\mathbf{Z}_{l};V)),$ (39)
Therefore, an outer bound for a given $(k,l)$ in Definition 3 is
$\mathcal{O}_{G_{kl}}~{}{\triangleq}$ $\bigcup_{P_{V|U},P_{U|\tilde{\mathbf{X}}}}\{(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:~{}$
$R_{S}$ $\leq I(\mathbf{Y}_{k};U|V)-I(\mathbf{Z}_{l};U|V),~{}R_{J}\geq I(\tilde{\mathbf{X}};U|\mathbf{Y}_{k}),$
$R_{L}$ $\geq I(X;U|\mathbf{Y}_{k})+I(\mathbf{Y}_{k};V)-I(\mathbf{Z}_{l};V)\},$ (40)
where $V$ and $U$ satisfy $V-U-\tilde{\mathbf{X}}-X-(\mathbf{Y}_{k},\mathbf{Z}_{l})$. Hence,
$\mathcal{R}_{G}\subseteq$ $\bigcap_{k\in\mathcal{K}}\bigcap_{l\in\mathcal{L}}\mathcal{O}_{G_{kl}}.$ (41)
The proof for the CS model can be derived using the same reasoning using [20, Th. 4] instead of [20, Th. 3].
Appendix B Proof of Lemma 1
Denote the covariance matrix of $(X,\mathbf{Y}_{k},\mathbf{Z}_{l})$ as ${\bf\Sigma}$, where
${\bf\Sigma}=\begin{bmatrix}\sigma^{2}_{X}&{\bf\Sigma}_{X\mathbf{Y}_{k}}&{\bf\Sigma}_{X\mathbf{Z}_{l}}\\ {\bf\Sigma}_{\mathbf{Y}_{k}X}&{\bf\Sigma}_{\mathbf{Y}_{k}}&{\bf\Sigma}_{\mathbf{Y}_{k}\mathbf{Z}_{l}}\\ {\bf\Sigma}_{\mathbf{Z}_{l}X}&{\bf\Sigma}_{\mathbf{Z}_{l}\mathbf{Y}_{k}}&{\bf\Sigma}_{\mathbf{Z}_{l}}\end{bmatrix}.$ (42)
Note that (2) depends on the marginal distribution of $(X,\mathbf{Y}_{k})$, (3) depends on the marginal distribution of $X$, and (5) and (6) depend on the marginal distribution of $(X,\mathbf{Z}_{l})$. Therefore, without loss of generality, using [45, Th. 3.5.2] and (42), it suffices to consider
$\mathbf{Y}_{k}$ $={\bf\Sigma}_{\mathbf{Y}_{k}X}\sigma^{-2}_{X}X+\mathbf{N}_{\mathbf{Y}_{k}},$ (43)
$\mathbf{Z}_{l}$ $={\bf\Sigma}_{\mathbf{Z}_{l}X}\sigma^{-2}_{X}X+\mathbf{N}_{\mathbf{Z}_{l}},$ (44)
where $\mathbf{N}_{\mathbf{Y}_{k}}\sim\mathcal{N}(0,{\bf\Sigma}_{\mathbf{N}_{\mathbf{Y}_{k}}})$ with ${\bf\Sigma}_{\mathbf{N}_{\mathbf{Y}_{k}}}={\bf\Sigma}_{\mathbf{Y}_{k}}-{\bf\Sigma}_{\mathbf{Y}_{k}X}\sigma^{-2}_{X}{\bf\Sigma}_{X\mathbf{Y}_{k}}$, and $\mathbf{N}_{\mathbf{Z}_{l}}\sim\mathcal{N}(0,{\bf\Sigma}_{\mathbf{N}_{\mathbf{Z}_{l}}})$ with ${\bf\Sigma}_{\mathbf{N}_{\mathbf{Z}_{l}}}={\bf\Sigma}_{\tilde{\mathbf{X}}}-{\bf\Sigma}_{\mathbf{Z}_{l}X}\sigma^{-2}_{X}{\bf\Sigma}_{X\mathbf{Z}_{l}}$, independent of $X$.
Since ${\bf\Sigma}$ is non-singular (positive definite), the sub-matrix $\begin{bmatrix}\sigma^{2}_{X}&{\bf\Sigma}_{X\mathbf{Y}_{k}}\\ {\bf\Sigma}_{\mathbf{Y}_{k}X}&{\bf\Sigma}_{\mathbf{Y}_{k}}\end{bmatrix}$ is also positive definite. This implies that the matrix ${\bf\Sigma}_{\mathbf{N}_{\mathbf{Y}_{k}}}$ is positive definite as it is the Schur complement of $\sigma^{2}_{X}$ in the sub-matrix. By Cholesky decomposition, there exists an invertible matrix ${\bf C}\in\mathbb{R}^{\Omega_{Y}\times\Omega_{Y}}$ such that ${\bf\Sigma}_{\mathbf{N}_{\mathbf{Y}_{k}}}={\bf C}{\bf C}^{\intercal}$. Then, we can reformulate (43) as
$\mathbf{Y}^{\prime}_{k}={\bf A}_{\mathbf{Y}^{\prime}_{k}}X+\mathbf{N}^{\prime}_{\mathbf{Y}_{k}},$ (45)
where $\mathbf{Y}^{\prime}_{k}={\bf C}^{-1}\mathbf{Y}_{k}$, ${\bf A}_{\mathbf{Y}^{\prime}_{k}}={\bf C}^{-1}{\bf\Sigma}_{\mathbf{Y}_{k}X}\sigma^{-2}_{X}$ and $\mathbf{N}^{\prime}_{\mathbf{Y}_{k}}\sim\mathcal{N}({\bf 0},{\bf I}_{\Omega_{Y}})$. Similarly, the same approach can be applied to (44).
Appendix C Proof of Theorem 1
This appendix consists of two parts, that is, the achievability part in Appendix C-A and the converse part in Appendix C-B.
C-A Achievability Proof
Note that Proposition 1 was proved under finite source alphabets. However, the result can be extended to Gaussian sources as well by employing a fine quantization before encoding and decoding processes, similar to [46].
By choosing $V$ as a constant in Proposition 1, the following regions are achievable.
$R_{G}~{}\supseteq~{}\bigcup_{P_{U|X}}$$\Big{\{}(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:$
$R_{S}$ $\leq\min_{k}I(\mathbf{Y}_{k};U)-\max_{l}I(\mathbf{Z}_{l};U),$
$R_{J}$ $\geq I(X;U)-\min_{k}I(\mathbf{Y}_{k};U),$
$R_{L}$ $\geq I(X;U)-\min_{l}I(\mathbf{Y}_{k};U)\Big{\}},$ (46)
$R_{C}~{}\supseteq~{}\bigcup_{P_{U|X}}$$\Big{\{}(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:$
$R_{S}$ $\leq\min_{k}I(\mathbf{Y}_{k};U)-\max_{l}I(\mathbf{Z}_{l};U),$
$R_{J}$ $\geq I(X;U)-\max_{l}I(\mathbf{Z}_{l};U),$
$R_{L}$ $\geq I(X;U)-\min_{l}I(\mathbf{Y}_{k};U)\Big{\}},$ (47)
where auxiliary random variable $U$ satisfies the Markov chain $U-X-(\mathbf{Y}_{k},\mathbf{Z}_{l})$ for all $k\in\mathcal{K}$ and $l\in\mathcal{L}$.
Lemma 4.
(Weinstein–Aronszajn Identity [47, Appx. B]) For any $a\in\mathbb{R}_{+}$ and matrix $\mathbf{H}\in\mathbb{R}^{\Omega\times 1}$, we have
${\det}(\mathbf{H}a\mathbf{H}^{\intercal}+{\bf I}_{\Omega})=a\mathbf{H}^{\intercal}\mathbf{H}+1,$
where ${\det}(\cdot)$ denotes the determinant of a matrix.
Lemma 5.
For given $\alpha\in(0,1]$, the function
$f(\mathbf{H}^{\intercal}\mathbf{H})=\log\left(\frac{\sigma^{2}_{X}\mathbf{H}^{\intercal}\mathbf{H}+1}{\alpha\sigma^{2}_{X}\mathbf{H}^{\intercal}\mathbf{H}+1}\right)$
is monotonically increasing with respect to $\mathbf{H}^{\intercal}\mathbf{H}$.
For $0<\alpha\leq 1$, consider
$X~{}{\triangleq}~{}U+\Theta,$ (48)
where $U\sim\mathcal{N}(0,(1-\alpha)\sigma^{2}_{X}))$ and $\Theta\sim\mathcal{N}(0,\alpha\sigma^{2}_{X})$. This relation implies that
$I(X;U)=\frac{1}{2}\log\left(\frac{1}{\alpha}\right).$ (49)
From (16) and (48), it follows that
$\mathbf{Y}_{k}$ $=\mathbf{H}_{k}U+\mathbf{H}_{k}\Theta+\mathbf{N}_{\mathbf{Y}_{k}},$
$\mathbf{Z}_{l}$ $=\tilde{\mathbf{H}}_{l}U+\tilde{\mathbf{H}}_{l}\Theta+\mathbf{N}_{\mathbf{Z}_{l}}.$
Using Lemma 4, we have
$I(\mathbf{Y}_{k};U)$ $=\frac{1}{2}\log\frac{\sigma^{2}_{X}\mathbf{H}_{k}^{\intercal}\mathbf{H}_{k}+1}{\alpha\sigma^{2}_{X}\mathbf{H}_{k}^{\intercal}\mathbf{H}_{k}+1},$
$I(\mathbf{Z}_{l};U)$ $=\frac{1}{2}\log\frac{\sigma^{2}_{X}\tilde{\mathbf{H}}^{\intercal}_{l}\tilde{\mathbf{H}}_{l}+1}{\alpha\sigma^{2}_{X}\tilde{\mathbf{H}}^{\intercal}_{l}\tilde{\mathbf{H}}_{l}+1}$
for a fixed pair $(k,l)$, and invoking Lemma 5 gives
$\min_{k}I(\mathbf{Y}_{k};U)$ $=\frac{1}{2}\log\left(\frac{\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}+1}{\alpha\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}+1}\right),$ (50)
$\max_{l}I(\mathbf{Z}_{l};U)$ $=\frac{1}{2}\log\left(\frac{\sigma^{2}_{X}\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\tilde{\mathbf{H}}_{l^{*}}+1}{\alpha\sigma^{2}_{X}\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\tilde{\mathbf{H}}_{l^{*}}+1}\right).$ (51)
Finally, substituting (49), (50), (51) into (46) and (47), gives (22) and (23).
C-B Converse Proof
We will need the following lemmas. These lemmas convert vector observations in (16) into scalar Gaussian random variables using sufficient statistics [39, Sect. 2.9]. This transformation plays an important role in deriving the outer bound of Gaussian sources.
Lemma 6.
([48, Lemma 3.1]) Consider a channel with input $W$ and output $\tilde{\bf W}$, namely, $\tilde{\bf W}~{}{\triangleq}~{}{\bf A}W+{\bf N}_{\tilde{\bf W}},$ where ${\bf N}_{\tilde{\bf W}}\sim\mathcal{N}({\bf 0},{\bf\Sigma}_{\tilde{\bf W}})$. A sufficient statistic to correctly determine $W$ from $\tilde{\bf W}$ is the following scalar
$\bar{W}~{}{\triangleq}~{}{\bf A}^{\intercal}{\bf\Sigma}^{-1}_{\tilde{\bf W}}\tilde{\bf W}.$ (52)
Lemma 7.
The vector equations in (16) can be rewritten as
$\bar{Y}_{k}=\nu_{\bar{Y}_{k}}X+N_{\bar{Y}_{k}},\bar{Z}_{l}=\nu_{\bar{Z}_{l}}X+N_{\bar{Z}_{l}},$ (53)
where $\nu_{\bar{Y}_{k}}\triangleq\mathbf{H}^{\intercal}_{k}\mathbf{H}_{k}$, $\nu_{\bar{Z}_{l}}\triangleq\tilde{\mathbf{H}}^{\intercal}_{l}\tilde{\mathbf{H}}_{l}$, $N_{\bar{Y}_{k}}\sim\mathcal{N}(0,\nu_{\bar{Y}_{k}})$, and $N_{\bar{Z}_{l}}\sim\mathcal{N}(0,\nu_{\bar{Z}_{l}})$.
Applying Lemma 6 to our settings in (16), we have
$\bar{Y}_{k}~{}=~{}\mathbf{H}^{\intercal}_{k}{\bf I}^{-1}_{\Omega_{Y}}\mathbf{Y}_{k},~{}\bar{Z}_{l}~{}=~{}\tilde{\mathbf{H}}^{\intercal}_{l}{\bf I}^{-1}_{\Omega_{Z}}\mathbf{Z}_{l}.$ (54)
Now substituting (16) into (54), we have
$\bar{Y}_{k}$ $=\mathbf{H}^{\intercal}_{k}(\mathbf{H}_{k}X+\mathbf{N}_{\mathbf{Y}_{k}})=\nu_{\bar{Y}_{k}}X+N_{\bar{Y}_{k}},$ (55)
$\bar{Z}_{l}$ $=\tilde{\mathbf{H}}^{\intercal}_{l}(\tilde{\mathbf{H}}_{l}X+\mathbf{N}_{\mathbf{Z}_{l}})=\nu_{\bar{Z}_{l}}X+N_{\bar{Z}_{l}},$ (56)
where we denote $\nu_{\bar{Y}_{k}}~{}{\triangleq}~{}\mathbf{H}^{\intercal}_{k}\mathbf{H}_{k}$, $\nu_{\bar{Z}_{l}}\triangleq\tilde{\mathbf{H}}^{\intercal}_{l}\tilde{\mathbf{H}}_{l}$, $N_{\bar{Y}_{k}}~{}{\triangleq}~{}\mathbf{H}_{k}\mathbf{N}_{\mathbf{Y}_{k}}$, and $N_{\bar{Z}_{l}}~{}{\triangleq}~{}\tilde{\mathbf{H}}_{l}\mathbf{N}_{\mathbf{Z}_{l}}$. Note that $N_{\bar{Y}_{k}}$ and $N_{\bar{Z}_{l}}$ are Gaussian random variables, and their variances are ${\rm Var}[\mathbf{H}^{\intercal}_{k}\mathbf{N}_{\mathbf{Y}_{k}}]=\mathbf{H}^{\intercal}_{k}\mathbf{H}_{k}$, and ${\rm Var}[\tilde{\mathbf{H}}^{\intercal}_{l}\mathbf{N}_{\mathbf{Z}_{l}}]=\tilde{\mathbf{H}}^{\intercal}_{l}\tilde{\mathbf{H}}_{l}$. ∎
As $(X,\bar{Y}_{k},\bar{Z}_{l})$ are scalar Gaussian random variables, when the squared value of the correlation coefficient of $(X,\bar{Y}_{k^{*}})$ is greater than that of $(X,\bar{Z}_{l^{*}})$, i.e., $\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}\geq\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\tilde{\mathbf{H}}_{l^{*}}$, implying that $\mathbf{H}^{\intercal}_{k}\mathbf{H}_{k}\geq\tilde{\mathbf{H}}^{\intercal}_{l}\mathbf{H}_{l}$ for any pair $(k,l)$, there exist Gaussian random variables $(X^{\prime},\bar{Y}^{\prime}_{k},\bar{Z}^{\prime}_{l})$ such that $X^{\prime}-\bar{Y}^{\prime}_{k}-\bar{Z}^{\prime}_{l}$ is satisfied, and the marginal distributions of $(X,\bar{Y}_{k})$ and $(X^{\prime},\bar{Y}^{\prime}_{k})$ and that of $(X,\bar{Z}_{l})$ and $(X^{\prime},\bar{Z}^{\prime}_{l})$ coincide [40, Lemma 6]. In the subsequent discussions, we denote $(X^{\prime},\bar{Y}^{\prime}_{k},\bar{Z}^{\prime}_{l})$ as $(X,\bar{Y}_{k},\bar{Z}_{l})$ for brevity. This property is used in the derivation of the following lemma.
Lemma 8 (Outer bounds).
If $\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}\geq\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\mathbf{H}_{l^{*}}$, the outer bounds of the GS and CS models are provided as
$\mathcal{R}_{G}~{}$ $\subseteq~{}\bigcap_{k\in\mathcal{K}}\bigcap_{l\in\mathcal{L}}\bar{\mathcal{O}}_{G_{kl}},~{}~{}~{}\mathcal{R}_{C}\subseteq~{}\bigcap_{k\in\mathcal{K}}\bigcap_{l\in\mathcal{L}}\bar{\mathcal{O}}_{C_{kl}},$ (57)
where $\bar{\mathcal{O}}_{G_{kl}}$ and $\bar{\mathcal{O}}_{C_{kl}}$ are outer bounds of the GS and CS models for a given pair $(k,l)$ and are defined as
$\bar{\mathcal{O}}_{G_{kl}}\triangleq\bigcup_{P_{U|X}}$$\Big{\{}(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:~{}$
$R_{S}$ $\leq I(\bar{Y}_{k};U)-I(\bar{Z}_{l};U),$
$R_{J}$ $\geq I(X;U)-I(\bar{Y}_{k};U),$
$R_{L}$ $\geq I(X;U)-I(\bar{Y}_{k};U)\Big{\}},$ (58)
$\bar{\mathcal{O}}_{C_{kl}}\triangleq\bigcup_{P_{U|X}}$$\Big{\{}(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:~{}$
$R_{S}$ $\leq I(\bar{Y}_{k};U)-I(\bar{Z}_{l};U),$
$R_{J}$ $\geq I(X;U)-I(\bar{Z}_{l};U),$
$R_{L}$ $\geq I(X;U)-I(\bar{Y}_{k};U)\Big{\}},$ (59)
and $U$ satisfies the Markov chain $U-X-\bar{Y}_{k}-\bar{Z}_{l}$. If $\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}<\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\mathbf{H}_{l^{*}}$, $\mathcal{R}_{G}=\mathcal{R}_{C}=\{(R_{S},R_{J},R_{L}):R_{S}=0,R_{J}\geq 0,R_{L}\geq 0\}.$
Proof.
See Appendix D. ∎
Next, we utilize the single-letter expressions in Lemma 8 to derive the parametric forms for Gaussian sources. We begin with the proof of (58). Each rate constraint in (58) can be expanded as
$R_{S}$ $\leq I(\bar{Y}_{k};U)-I(\bar{Z}_{l};U)\overset{\rm(a)}{=}\frac{1}{2}\log\frac{\mathbf{H}^{\intercal}_{k}\mathbf{H}_{k}(\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k}\mathbf{H}_{k}+1)}{\tilde{\mathbf{H}}^{\intercal}_{l}\tilde{\mathbf{H}}_{l}(\sigma^{2}_{X}\mathbf{H}^{\intercal}_{l}\mathbf{H}_{l}+1)}$
$~{}~{}~{}~{}+h(\bar{Z}_{l}|U)-h(\bar{Y}_{k}|U),$ (60)
$R_{J}$ $\geq I(X;U)-I(\bar{Y}_{k};U)\overset{\rm(b)}{=}\frac{1}{2}\log\frac{\sigma^{2}_{X}}{\mathbf{H}^{\intercal}_{k}\mathbf{H}_{k}(\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k}\mathbf{H}_{k}+1)}$
$~{}~{}~{}~{}+h(\bar{Y}_{k}|U)-h(X|U),$ (61)
$R_{L}$ $\geq I(X;U)-I(\bar{Y}_{k};U)\overset{\rm(c)}{=}\frac{1}{2}\log\frac{\sigma^{2}_{X}}{\mathbf{H}^{\intercal}_{k}\mathbf{H}_{k}(\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k}\mathbf{H}_{k}+1)}$
$~{}~{}~{}~{}+h(\bar{Y}_{k}|U)-h(X|U),$ (62)
where (a), (b), and (c) follow from (53).
From Lemma 7, we have
$\frac{1}{2}$ $\log\frac{\sigma^{2}_{X}+1/\nu_{\bar{Z}_{l}}}{\sigma^{2}_{X}+1/\nu_{\bar{Y}_{k}}}=h(\frac{1}{\nu_{\bar{Z}_{l}}}\bar{Z}_{l})-h(\frac{1}{\nu_{\bar{Y}_{k}}}\bar{Y}_{k})$
$\overset{\rm(a)}{\leq}h(\frac{1}{\nu_{\bar{Z}_{l}}}\bar{Z}_{l}|U)-h(\frac{1}{\nu_{\bar{Y}_{k}}}\bar{Y}_{k}|U)$
$\overset{\rm(b)}{\leq}h(\frac{1}{\nu_{\bar{Z}_{l}}}\bar{Z}_{l}|X)-h(\frac{1}{\nu_{\bar{Y}_{k}}}\bar{Y}_{k}|X)=\frac{1}{2}\log\frac{1/\nu_{\bar{Z}_{l}}}{1/\nu_{\bar{Y}_{k}}},$ (63)
where (a) and (b) follow from the fact that $I(\bar{Y}_{k};U|\bar{Z}_{l})\geq 0$ and $I(\bar{Y}_{k};X|U,\bar{Z}_{l})\geq 0$, respectively. Thus, there must exist a parameter $\alpha\in(0,1]$ such that
$h(\frac{1}{\nu_{\bar{Z}_{l}}}\bar{Z}_{l}|U)-h(\frac{1}{\nu_{\bar{Y}_{k}}}\bar{Y}_{k}|U)=\frac{1}{2}\log\frac{\alpha\sigma^{2}_{X}+1/\nu_{\bar{Z}_{l}}}{\alpha\sigma^{2}_{X}+1/\nu_{\bar{Y}_{k}}}.$ (64)
Equation (64) also indicates that
$h(\bar{Z}_{l}|U)-h(\bar{Y}_{k}|U)$ $=\frac{1}{2}\log\frac{\nu_{\bar{Z}_{l}}(\alpha\sigma^{2}_{X}\nu_{\bar{Z}_{l}}+1)}{\nu_{\bar{Y}_{k}}(\alpha\sigma^{2}_{X}\nu_{\bar{Y}_{k}}+1)}$
$=\frac{1}{2}\log\frac{\tilde{\mathbf{H}}^{\intercal}_{l}\tilde{\mathbf{H}}_{l}(\alpha\sigma^{2}_{X}\tilde{\mathbf{H}}^{\intercal}_{l}\tilde{\mathbf{H}}_{l}+1)}{\mathbf{H}^{\intercal}_{k}\mathbf{H}_{k}(\alpha\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k}\mathbf{H}_{k}+1)}.$ (65)
The conditional Fisher information of $A$ is defined by $\mathbb{J}(A|U)=\mathbb{E}\left[\left(\frac{\partial\log f_{A|U}(a|u)}{\partial a}\right)^{2}\right],$ where the expectation is taken over $(U,A)$ [42, Def. 1].
Lemma 9.
([42, Cor. 1]) Let $W,A,B$ be random variables, and let the density for any combination of them exist. Moreover, assume that given $W$, $A$ and $B$ are independent. Then, we have
$\frac{1}{\mathbb{J}(A+B|W)}\geq\frac{1}{\mathbb{J}(A|W)}+\frac{1}{\mathbb{J}(B|W)}.$ (66)
Lemma 10.
For $0\leq r\leq 1/\bar{\nu}_{\bar{Y}_{k}}$, it holds that
$\mathbb{J}(X+\sqrt{r}N|U)\geq\frac{1}{\alpha\sigma^{2}_{X}+r}$ (67)
with an independent Gaussian random variable $N\sim\mathcal{N}(0,1)$.
Proof.
From [42, Lemma 3], it follows that
[…]
with an independent random variable $\tilde{N}\sim\mathcal{N}(0,1)$. Then,
[…]
=$\overset{\rm(a)}{=}\frac{1}{2}\int_{1/\nu_{\bar{Y}_{k}}}^{1/\nu_{\bar{Z}_{l}}}\mathbb{J}(X+\sqrt{r}N+\sqrt{t-r}N^{\prime}|U)dt$
$\overset{\rm(b)}{\leq}\frac{1}{2}\int_{1/\nu_{\bar{Y}_{k}}}^{1/\nu_{\bar{Z}_{l}}}\left(\mathbb{J}(X+\sqrt{r}N|U)^{-1}+t-r\right)^{-1}dt$
[…]
$=\frac{1}{2}\log\frac{1+\mathbb{J}(X+\sqrt{r}N|U)(1/\nu_{\bar{Z}_{l}}-r)}{1+\mathbb{J}(X+\sqrt{r}N|U)(1/\nu_{\bar{Y}_{k}}-r)},$ (68)
where (a) follows by picking a real number $r$ in the range of $0\leq r\leq 1/\bar{\nu}_{\bar{Y}_{k}}$ and using independent Gaussian random variables $N\sim\mathcal{N}(0,1)$ and $N^{\prime}\sim\mathcal{N}(0,1)$, and (b) is due to Lemma 9. Lastly, comparing (64) and (68), we obtain (67). ∎
Observe that
[…]
$\overset{\rm(a)}{\geq}\frac{1}{2}\int_{0}^{1/\nu_{\bar{Y}_{k}}}\frac{1}{\alpha\sigma^{2}_{X}+r}dr=\frac{1}{2}\log\frac{\alpha\sigma^{2}_{X}+1/\nu_{\bar{Y}_{k}}}{\alpha\sigma^{2}_{X}},$ (69)
where (a) follows from Lemma 10, which implies that
$h(\bar{Y}_{k}|U)-h(X|U)$ $\geq\frac{1}{2}\log\frac{\nu_{\bar{Y}_{k}}(\alpha\sigma^{2}_{X}\nu_{\bar{Y}_{k}}+1)}{\alpha\sigma^{2}_{X}}$
$=\frac{1}{2}\log\frac{\mathbf{H}^{\intercal}_{k}\mathbf{H}_{k}(\alpha\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k}\mathbf{H}_{k}+1)}{\alpha\sigma^{2}_{X}}.$ (70)
Substituting (65) and (70) into the rate constraints in (60)–(62), the outer bound for a fixed pair $(k,l)$ is expressed as
$\bar{\mathcal{O}}_{G_{kl}}$ $\triangleq\bigcup_{0<\alpha\leq 1}\Big{\{}(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:~{}$
$R_{S}$ $\leq\frac{1}{2}\log\frac{\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k}\mathbf{H}_{k}+1}{\alpha\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k}\mathbf{H}_{k}+1}-\frac{1}{2}\log\frac{\sigma^{2}_{X}\tilde{\mathbf{H}}^{\intercal}_{l}\tilde{\mathbf{H}}_{l}+1}{\alpha\sigma^{2}_{X}\tilde{\mathbf{H}}^{\intercal}_{l}\tilde{\mathbf{H}}_{l}+1},$
$R_{J}$ $\geq\frac{1}{2}\log\frac{\alpha\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k}\mathbf{H}_{k}+1}{\alpha(\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k}\mathbf{H}_{k}+1)},$
$R_{L}$ $\geq\frac{1}{2}\log\frac{\alpha\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k}\mathbf{H}_{k}+1}{\alpha(\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k}\mathbf{H}_{k}+1)}\Big{\}}.$ (71)
Applying Lemma 5 to (57) and (71), the outer region of the GS model for all possible index pairs $(k,l)$ is
$\mathcal{R}_{G}$ $\subseteq\bigcup_{0<\alpha\leq 1}\Big{\{}(R_{S},R_{J},R_{L})\in\mathbb{R}^{3}_{+}:~{}$
$R_{S}$ $\leq\frac{1}{2}\log\frac{(\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}+1)(\alpha\sigma^{2}_{X}\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\tilde{\mathbf{H}}_{l^{*}}+1)}{(\alpha\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}+1)(\sigma^{2}_{X}\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\tilde{\mathbf{H}}_{l^{*}}+1)},$
$R_{J}$ $\geq\frac{1}{2}\log\frac{\alpha\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}+1}{\alpha(\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}+1)},$
$R_{L}$ $\geq\frac{1}{2}\log\frac{\alpha\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}+1}{\alpha(\sigma^{2}_{X}\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}+1)}\Big{\}}.$ (72)
The outer region of the CS model in (23) can be shown similarly. ∎
Appendix D Proof of Lemma 8
We only prove (58), the outer bound of the GS model for a given pair $(k,l)$, as the proof of (8) follows by a similar manner. Assume that a rate tuple $(R_{S},R_{J},R_{L})$ is achievable with respect to Definition 3 for every pair $(k,l)\in\mathcal{K}\times\mathcal{L}$.
We begin by establishing the following Markov chains:
$(J,S)-X^{n}-\mathbf{Y}^{n}_{k}-\bar{Y}^{n}_{k},~{}(J,S)-X^{n}-\bar{Y}^{n}_{k}-\mathbf{Y}^{n}_{k},$ (73)
$(J,S)-X^{n}-\mathbf{Z}^{n}_{l}-\bar{Z}^{n}_{l},~{}~{}(J,S)-X^{n}-\bar{Z}^{n}_{l}-\mathbf{Z}^{n}_{l},$ (74)
$(J,S)-X^{n}-(\mathbf{Y}^{n}_{k},\mathbf{Z}^{n}_{l})-(\bar{Y}^{n}_{k},\bar{Z}^{n}_{l}),$ (75)
where the left-hand sides of (73) and (74), and (75) hold because $\bar{Y}^{n}_{k}$ and $\bar{Z}^{n}_{l}$ are functions of $\mathbf{Y}^{n}_{k}$ and $\mathbf{Z}^{n}_{l}$, respectively, by Lemma 7, and the right-hand sides of (73) and (74) are due to the sufficient statistic [39, Sect. 2.9]. In addition, for the scalar random variables $(X,\bar{Y}_{k},\bar{Z}_{l})$, the Markov chain $X^{n}-\bar{Y}^{n}_{k}-\bar{Z}^{n}_{l}$ holds for any pair $(k,l)$ when $\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}\geq\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\tilde{\mathbf{H}}_{l^{*}}$. Combining this with (75) gives
$(J,S)-X^{n}-\bar{Y}^{n}_{k}-\bar{Z}^{n}_{l}.$ (76)
Define auxiliary random variables
$V_{t}=(J,\bar{Y}^{n}_{k,t+1},\bar{Z}^{t-1}_{l})~{}{\rm and}~{}U_{t}=(J,S,\bar{Y}^{n}_{k,t+1},\bar{Z}^{t-1}_{l}),$ (77)
which guarantee the Markov chain
$V_{t}-U_{t}-X_{t}-\bar{Y}_{k,t}-\bar{Z}_{l,t}.$ (78)
Also, we define
$\delta_{n}=\frac{1}{n}\left(H_{b}(\delta)+\delta\log|\mathcal{S}|\right),$ (79)
where $H_{b}(\cdot)$ denotes the binary entropy function, and $\delta_{n}\downarrow 0$ as $\delta\downarrow 0$ and $n\rightarrow\infty$.
Analysis of Secret-Key Rate: From (3),
$n(R_{S}-\delta)\leq H(S)$
$=H(S|J,\mathbf{Z}^{n}_{l})+I(S;J,\mathbf{Z}^{n}_{l})$
$\overset{\rm(a)}{\leq}H(S|J,\mathbf{Z}^{n}_{l})-H(S|J,\mathbf{Y}^{n}_{k})+n(\delta+\delta_{n})$
=$\overset{\rm(b)}{=}H(S|J,\mathbf{Z}^{n}_{l},\bar{Z}^{n}_{l})-H(S|J,\mathbf{Y}^{n}_{k},\bar{Y}^{n}_{k})+n(\delta+\delta_{n})$
=$\overset{\rm(c)}{=}H(S|J,\bar{Z}^{n}_{l})-H(S|J,\bar{Y}^{n}_{k})+n(\delta+\delta_{n})$
$=I(S;\bar{Y}^{n}_{k}|J)-I(S;\bar{Z}^{n}_{l}|J)+n(\delta+\delta_{n})$
=$\overset{\rm(d)}{=}\sum_{t=1}^{n}\{I(\bar{Y}_{k,t};U_{t}|V_{t})-I(\bar{Z}_{l,t};U_{t}|V_{t})\}+n(\delta+\delta_{n})$ (80)
=$\overset{\rm(e)}{=}\sum_{t=1}^{n}\{I(\bar{Y}_{k,t};U_{t})-I(\bar{Z}_{l,t};U_{t})$
$~{}~{}~{}-(I(\bar{Y}_{k,t};V_{t})-I(\bar{Z}_{l,t};V_{t}))\}+n(\delta+\delta_{n})$
$\overset{\rm(f)}{\leq}\sum_{t=1}^{n}\{I(\bar{Y}_{k,t};U_{t})-I(\bar{Z}_{l,t};U_{t})\}+n(\delta+\delta_{n}),$ (81)
where (a) is due to (5) and Fano’s inequality with $\delta_{n}$ defined in (79) as the secret key $S$ can be reliably estimated from $(J,\mathbf{Y}^{n}_{k})$, (b) follows from the left-hand sides of (73) and (74), (c) holds by the right-hand sides of (73) and (74), (d) follows by [49, Lemma 4.1], (e) follows from the Markov chains $V_{t}-U_{t}-\bar{Y}_{k,t}$ and $V_{t}-U_{t}-\bar{Z}_{l,t}$, and (f) is due to (78), which results in $I(\bar{Y}_{k,t};V_{t})-I(\bar{Z}_{l,t};V_{t})\geq 0$.
Analysis of Storage Rate: From (4),
$n$ $(R_{J}+\delta)\geq\log|\mathcal{J}|\geq H(J)=I(X^{n};J)$
=$\overset{\rm(a)}{=}I(X^{n};J|\mathbf{Y}^{n}_{k})$
$\geq I(X^{n};J,S|\mathbf{Y}^{n}_{k})-H(S|\mathbf{Y}^{n}_{k},J)$
$\overset{\rm(b)}{\geq}I(X^{n};J,S|\mathbf{Y}^{n}_{k})-n\delta_{n}$
$=I(X^{n};\bar{Y}^{n}_{k}|\mathbf{Y}^{n}_{k})+I(X^{n};J,S|\bar{Y}^{n}_{k},\mathbf{Y}^{n})$
$~{}~{}~{}~{}-I(X^{n};\bar{Y}^{n}_{k}|J,S,\mathbf{Y}^{n}_{k})-n\delta_{n}$
=$\overset{\rm(c)}{=}I(X^{n};J,S|\bar{Y}^{n}_{k},\mathbf{Y}^{n}_{k})-n\delta_{n}$
=$\overset{\rm(d)}{=}I(X^{n};J,S|\bar{Y}^{n}_{k})-n\delta_{n}$ (82)
=$\overset{\rm(f)}{=}\sum_{t=1}^{n}\{h(X_{t}|\bar{Y}_{k,t})-h(X_{t}|J,S,X^{t-1},\bar{Y}^{n}_{k},\bar{Z}^{t-1}_{l})\}-n\delta_{n}$
$\overset{\rm(f)}{\geq}\sum_{t=1}^{n}\{h(X_{t}|\bar{Y}_{k,t})-h(X_{t}|U_{t},\bar{Y}_{k,t})\}-n\delta_{n}$
=$\overset{\rm(g)}{=}\sum_{t=1}^{n}\{I(X_{t};U_{t})-I(\bar{Y}_{k,t};U_{t})\}-n\delta_{n},$ (83)
where (a) is due to the Markov chain $J-X^{n}-\mathbf{Y}^{n}_{k}$, (b) follows by Fano’s inequality with $\delta_{n}$ defined in (79), (c) follows because $\bar{Y}^{n}_{k}$ is a function of $\mathbf{Y}^{n}_{k}$, (d) holds from the left-hand side of (73), (e) holds due to the Markov chain $X_{t}-(J,S,X^{t-1},\bar{Y}^{n}_{k},)-\bar{Z}^{t-1}_{l}$, (f) follows because conditioning reduces entropy, and (g) is due to $U_{t}-X_{t}-\bar{Y}_{k,t}$.
Analysis of Privacy-Leakage Rate: For a fixed $l$, we first show that the left-hand side of (6) is preserved when $(X^{n},\mathbf{Z}^{n})$ is replaced with $(X^{n},\bar{Z}^{n})$.
$I(X^{n};J|\mathbf{Z}^{n}_{l})$ =$\overset{\rm(a)}{=}I(X^{n};J)-I(\mathbf{Z}^{n}_{l};J)$
=$\overset{\rm(b)}{=}I(X^{n};J)-I(\mathbf{Z}^{n}_{l},\bar{Z}^{n}_{l};J)$
=$\overset{\rm(c)}{=}I(X^{n};J)-I(\bar{Z}^{n}_{l};J)$
=$\overset{\rm(d)}{=}I(X^{n};J|\bar{Z}^{n}_{l}),$ (84)
where (a), (b), (c), and (d) follow from the Markov chains $J-X^{n}-\mathbf{Z}^{n}_{l}$, $J-\mathbf{Z}^{n}_{l}-\bar{Z}^{n}_{l}$, $J-\bar{Z}^{n}_{l}-\mathbf{Z}^{n}_{l}$, and $J-X^{n}_{l}-\bar{Z}^{n}_{l}$, respectively, all of which are obtained as special cases of (74).
Therefore, we can evaluate the privacy-leakage rate as
$n(R_{L}+\delta)\geq I(X^{n};J|\mathbf{Z}^{n}_{l})$
=$\overset{\rm(a)}{=}I(X^{n};J|\bar{Z}^{n}_{l})$
$=I(X^{n};J,S,\mathbf{Y}^{n}_{k}|\bar{Z}^{n}_{l})-I(X^{n};\mathbf{Y}^{n}_{k}|J,\bar{Z}^{n}_{l})$
$~{}~{}-I(X^{n};S|J,\mathbf{Y}^{n}_{k},\bar{Z}^{n}_{l})$
$\geq I(X^{n};J,S,\mathbf{Y}^{n}_{k}|\bar{Z}^{n}_{l})-I(X^{n};\mathbf{Y}^{n}_{k}|J,\bar{Z}^{n}_{l})-H(S|J,\mathbf{Y}^{n}_{k})$
$\geq I(X^{n};J,S,\mathbf{Y}^{n}_{k}|\bar{Z}^{n}_{l})-I(X^{n};\mathbf{Y}^{n}_{k}|J,\bar{Z}^{n}_{l})-n\delta_{n}$
=$\overset{\rm(b)}{=}I(X^{n};J,S,\bar{Y}^{n}_{k},\mathbf{Y}^{n}_{k}|\bar{Z}^{n}_{l})-I(X^{n};\bar{Y}^{n}_{k},\mathbf{Y}^{n}_{k}|J,\bar{Z}^{n}_{l})-n\delta_{n}$
=$\overset{\rm(c)}{=}I(X^{n};J,S,\bar{Y}^{n}_{k}|\bar{Z}^{n}_{l})-I(X^{n};\bar{Y}^{n}_{k}|J,\bar{Z}^{n}_{l})-n\delta_{n}$
=$\overset{\rm(d)}{=}I(X^{n};J,S|\bar{Y}^{n}_{k},\bar{Z}^{n}_{l})+I(X^{n};\bar{Y}^{n}_{k}|\bar{Z}^{n}_{l})$
$~{}~{}~{}-(I(X^{n};\bar{Y}^{n}_{k}|\bar{Z}^{n}_{l})-I(\bar{Y}^{n}_{k};J|\bar{Z}^{n}_{l}))-n\delta_{n}$
$\geq I(X^{n};J,S|\bar{Y}^{n}_{k},\bar{Z}^{n}_{l})-n\delta_{n}$
=$\overset{\rm(e)}{=}I(X^{n};J,S|\bar{Y}^{n}_{k})-n\delta_{n}$
$\overset{\rm(f)}{\geq}\sum_{t=1}^{n}\{I(X_{t};U_{t})-I(\bar{Y}_{k,t};U_{t})\}-n\delta_{n},$ (85)
where (a) is due to (84), (b) follows from (54), i.e., $\bar{Y}^{n}_{k}$ is a function of $\mathbf{Y}^{n}_{k}$, (c) holds because the Markov chain $X^{n}-(J,S,\bar{Y}^{n}_{k},\bar{Z}^{n}_{l})-\mathbf{Y}^{n}_{k}$ holds, from the right-hand side of (73), (d) and (e) are due to the Markov chain $(J,S)-X^{n}-Y^{n}_{k}-Z^{n}_{l}$, obtained from (76), and (e) follows by the same steps from (82) to (83).
For the case where $\mathbf{H}^{\intercal}_{k^{*}}\mathbf{H}_{k^{*}}<\tilde{\mathbf{H}}^{\intercal}_{l^{*}}\tilde{\mathbf{H}}_{l^{*}}$, the Markov chain $V_{t}-U_{t}-X_{t}-\bar{Z}_{l,t}-\bar{Y}_{k,t}$ holds. The secret-key rate follows from (80) since $I(\bar{Y}_{k,t};U_{t}|V_{t})\leq I(\bar{Z}_{l,t};U_{t}|V_{t})$.
Finally, we introduce a time-sharing random variable $Q\sim{\rm Unif}[1:n]$, independent of other random variables, and define $U=(U_{Q},Q)$, $X=X_{Q}$, $\bar{Y}_{k}=\bar{Y}_{k,Q}$, and $\bar{Z}_{l}=\bar{Z}_{l,Q}$, so that the Markov chain $U-X-\bar{Y}_{k}-\bar{Z}_{l}$ holds. By letting $n\rightarrow\infty$ and $\delta\downarrow 0$, one can see that for a given pair $(k,l)$, the outer bound of the GS model is given by (58). Hence, the outer bound valid for any pair $(k,l)$ is given by (57). ∎